### v1

# Let the Trial Begin: A Mock-Court Approach to Vulnerability Detection using LLM-Based Agents

Abstract.
Detecting vulnerabilities in source code remains a critical yet challenging task, especially when benign and vulnerable functions share significant similarities. In this work, we introduce VulTrial, a courtroom-inspired multi-agent framework designed to enhance automated vulnerability detection. It employs four role-specific agents, which are security researcher, code author, moderator, and review board. Through extensive experiments using GPT-3.5 and GPT-4o we demonstrate that VulTrial outperforms single-agent and multi-agent baselines. Using GPT-4o, VulTrial improves the performance by 102.39% and 84.17% over its respective baseline. Additionally, we show that role-specific instruction tuning in multi-agent with small data (50 pair samples) improves the performance of VulTrial further by 139.89% and 118.30%. Furthermore, we analyze the impact of increasing the number of agent interactions on VulTrial’s overall performance. While multi-agent setups inherently incur higher costs due to increased token usage, our findings reveal that applying VulTrial to a cost-effective model like GPT-3.5 can improve its performance by 69.89% compared to GPT-4o in a single-agent setting, at a lower overall cost.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Vulnerability detection in software code is a critical task to ensure system security and prevent potential exploits (Chen et al., 2023). Manual vulnerability detection, however, can be time-consuming and error-prone, especially given the scale and complexity of modern software systems. To automate this process, researchers have explored various machine learning and deep learning approaches (Zou et al., 2021a; Russell et al., 2018; Chernis and Verma, 2018; Risse and Böhme, 2024; Zou et al., 2021b; Chakraborty et al., 2021). Recently, with substantial advancements in large language models (LLMs) and their impressive performance on natural language tasks (Wang et al., 2023; Abbasiantaeb et al., 2024; Zhuang et al., 2023; Neumann et al., 2024), several studies have attempted to utilize LLMs for automated vulnerability detection tasks (Tamberg and Bahsi, 2024; Ullah et al., 2023; Zhou et al., 2024; Nong et al., 2024; Zhang et al., 2024; Widyasari et al., 2024), which shows promising outcomes.
Despite these advances, recent research by Ding et al. (Ding et al., 2025b) highlighted significant limitations when evaluating current LLM-based vulnerability detection approaches. Their study proposed a challenging new benchmark called PrimeVul, where the functional differences between benign and vulnerable code pairs are subtle, less than 20%. These cases where the code has a high degree of similarity are critical to evaluate because in real-world scenarios, vulnerabilities often stem from subtle, non-obvious code changes (Ni et al., 2023; Tan et al., 2024). On this dataset, LLM-based approaches exhibited low performance. The best performance comes from using GPT-4 (OpenAI, 2023b), which correctly predicts both codes in a pair (vulnerable and benign), referred to as Pair-wise Correct Prediction (P-C), 12.4% of cases. This LLM-based approach used in PrimeVul was implemented under the conventional single-agent system, wherein one model attempts to identify and explain vulnerabilities on its own.
Single-agent systems are inherently constrained by their reliance on a single perspective when analyzing vulnerabilities, which can lead to issues such as hallucination and overconfidence in incorrect reasoning (Zhang et al., 2023; Shi et al., 2025; Huang et al., 2025; Liang et al., 2023; Du et al., 2023). Moreover, these single-agent systems may fail to critically verify their own output, which is crucial when identifying subtle vulnerabilities (Liang et al., 2023).
Given these limitations, recent research has begun exploring multi-agent systems to enhance vulnerability detection (Wen et al., 2024; Hu et al., 2023; Wei et al., 2024). Multi-agent approaches distribute specific tasks among role-specific agents while allowing discussion and more information exchange between agents. GPTLens (Hu et al., 2023) is one such approach, dividing vulnerability detection into two stages, generation by an auditor agent and discrimination by a critic agent. In GPTLens, the final decision is made by the critic agent, which is tasked to be a harsh vulnerability critic and doubt the vulnerabilities and reasoning given by the auditor agent.
In this work, we propose VulTrial, a novel multi-agent LLM framework that employs a courtroom-inspired scenario. VulTrial features four role-specific agents: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). Unlike GPTLens, where the agent that does the critic delivers the final result, VulTrial delegates the final prediction to a dedicated review board agent, which is designed to be unbiased toward the security researcher and code author agents. The review board bases its decision on arguments from the security researcher (highlighting potential vulnerabilities), the code author (defending the code), and the moderator (summarizing debate points). Review board agent works similarly to the jury in the courtroom scenario, where they need to hear all the evidence presented in the case and make a final verdict based on it. Additionally, VulTrial allows for customizable rounds of interaction among the first three agents, which can provide opportunities for rebuttal and refinements.
We evaluate VulTrial results using two base LLMs, GPT-3.5 and GPT-4o (see Section 3.3). We benchmarked VulTrial against single-agent baselines from PrimeVul (Ding et al., 2025b) and a multi-agent baseline, GPTLens (Hu et al., 2023), all of which also utilize the same base LLMs for a fair comparison. PrimeVul (Ding et al., 2025b) (chain of thought and fine-tuning) and a multi-agent baseline, GPTLens (Hu et al., 2023). In the experiment using GPT-3.5, VulTrial outperforms the single-agent and multi-agent baselines by 151.69% and 239.78% respectively. Similarly, using GPT-4o, VulTrial also improves the performance of single-agent and multi-agent baselines by 102.39% and 84.17%.
Furthermore, to enhance VulTrial’s performance and analyze the impact of instruction tuning on a role-specific agent, we conducted additional experiments by applying instruction tuning to each agent in VulTrial. For the instruction tuning dataset, we selected 50 sample pairs from the validation data. The results indicate that tuning the moderator agent in VulTrial produced the most significant improvements, achieving an 18.53% increase compared to its non-tuned counterpart and outperforming the single-agent and multi-agent baselines by 139.89% and 118.30%, respectively. We also examined the effect of increasing the number of discussion rounds in VulTrial and observed that, when using GPT-4o, adding more rounds does not result in improved performance. Finally, an ablation study revealed that each agent within VulTrial is crucial for achieving the best overall performance. To summarize, the contributions of our study are as follows:
  • We propose VulTrial, a novel multi-agent, courtroom-inspired approach leveraging LLMs for automated vulnerability detection. VulTrial consists of role-specific agents (i.e., security researcher, code author, moderator, and review board). The ablation study highlights that each agent in VulTrial is necessary to get the best performance.
  • We demonstrate that targeted instruction tuning of role-specific agents in VulTrial improves vulnerability detection performance compared to the baselines. Even with limited training examples (as few as 50 pairs), we get an improvement of 18.53% compared to the non-tuned version.
  • We conduct extensive experiments using GPT-3.5 and GPT-4o in PrimeVul dataset (Ding et al., 2025b) and compare the VulTrial with single-agent and multi-agent baselines. Overall, the results using VulTrial outperform the baselines.
2. VulTrial
Our proposed method, namely VulTrial, works by utilizing multiple agents, all of which are powered by LLMs. VulTrial consists of four role-specific agents, each inspired by distinct roles in a courtroom setting: a security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). This structured debate encourages thorough, adversarial discussions, allowing for deeper analyses and refined conclusions regarding code vulnerabilities. Drawing inspiration from courtroom practices, where structured adversarial processes have been shown to reduce wrongful convictions and improve the fairness of judgments (Findley and Scott, 2006; Garrett, 2011), VulTrial aims to improve vulnerability detection in a similar manner. The overall architecture of these agents is illustrated in Fig. 1, with an example output in Fig. 2.
The first agent acts as a security researcher and is responsible to analyze the given code and identify potential vulnerabilities. The second agent, namely the code author agent, aims to provide counterarguments and justifications to refute the claims made by the security researcher agent. The third agent, which is the moderator agent, is tasked with summarizing the arguments presented by both the security researcher and the code author agents to ensure a structured exchange of viewpoints between them. VulTrial supports multiple rounds of interaction among the first three agents before passing their viewpoints into the last agent. The last agent, namely the review board agent, is tasked with analyzing all the information given by the moderator, the security researcher, and the code author agents. Based on this collective information, the review board delivers a final verdict regarding the security status of the code. This verdict highlights whether the potential vulnerability brought up by the security researcher agent is indeed a vulnerability. We chose to have the agents simulate a jury sentencing scenario over a judge sentencing scenario, as previous studies have shown that jury participation can help prevent excessively harsh punishments (Hans et al., 2015; McLeod, 2024). The details of each agent and their interaction will be discussed in the following subsections.
Figure 1. Overview of the VulTrial architecture. It employs a multi-agent framework inspired by a courtroom setting to assess code vulnerabilities. The security researcher (prosecutor) identifies potential vulnerabilities, the code author (defense attorney) provides counterarguments, the moderator (judge) summarizes key points, and the review board (jury) delivers the final verdict.
2.1. Security Researcher Agent (Prosecutor Role)
In VulTrial, the security researcher agent functions similarly to a prosecutor in a courtroom. A prosecutor’s role is to investigate evidence, establish the facts, and build a case against a defendant (i.e., the party that is accused in a court of law). In the case of vulnerability detection, the security researcher agent inspects the provided code, identifies potential vulnerabilities, and presents a rationale for why these vulnerabilities may pose a threat. An example of the security researcher agent’s output is shown in Fig. 2, where it highlights two potential vulnerabilities in the code, along with their reasoning and impact.
The prompt used for this agent is highlighted in Fig. 3. In the first round, the agent only receives a code function and determines whether the code is vulnerable. In subsequent rounds, it also receives the output generated by two other agents (i.e., code author and moderator) which is highlighted as ¡chat history¿ in Fig. 3. Beyond simply reporting any vulnerability, the agent is explicitly asked to include (1) a description of each vulnerability, (2) the reasoning behind its identification, and (3) the potential impact if that vulnerability were to be exploited. We asked for this agent to include the reasoning as it has been highlighted in previous studies (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024) that prompting LLMs to articulate their reasoning processes can significantly enhance their performance. Furthermore, including the potential impact of each vulnerability gives the other agents (i.e., the code author, moderator, and review board) a clearer perspective when they form their answer.
Figure 2. Example of vulnerable code (CVE-2021-37687) and analysis results from VulTrial. Green text indicates correct arguments, while orange text indicates misleading or exaggerated claims. The code is vulnerable as it does not check for negative indices, allowing an attacker to read arbitrary data by crafting a model with negative values in indices.
You are the Security Researcher.
Identify all potential security vulnerabilities in the given code snippet.
Provide your output as a JSON array. Each element in the array represents one identified vulnerability and should include:
vulnerability”: A short name or description of the vulnerability. reason”: A detailed explanation of why this is a vulnerability and how it could be exploited. impact”: The potential consequences if this vulnerability were exploited. ¡code¿
¡chat history¿
Figure 3. Security researcher agent prompt.
2.2. Code Author Agent (Defense Attorney Role)
The code author agent in VulTrial acts similarly to a defense attorney in a courtroom scenario. While the prosecutor aims to prove that the defendant is guilty, the defense attorney counters these allegations by advocating for the defendant’s innocence. Applying this scenario to the vulnerability detection, the code author agent seeks to defend the integrity and security of their code, providing arguments and evidence to counter the security researcher agent’s claims. If the vulnerability is valid, the agent can acknowledge the issue and propose effective mitigation strategies.
The specific prompt for the code author agent is shown in Fig. 4. This prompt provides the agent with the code snippet under investigation along with information generated by the security researcher agent (included in ¡chat history¿), which outlines a list of potential vulnerabilities. The agent is then tasked with clearly articulating the following outputs: (1) identification of the vulnerability being discussed, (2) explicit acknowledgment or refutation of the vulnerability, and (3) detailed reasoning justifying the refutation or, in cases where the agent acknowledges the vulnerability, proposed methods for mitigation. An example of the code author agent’s output is also shown in Fig. 2, where it acknowledges one of the vulnerabilities identified by the security researcher and proposes a mitigation strategy. Additionally, it refutes the other vulnerability and provides reasoning for it.
Since the security researcher agent may highlight multiple potential vulnerabilities, it is crucial for the code author agent to clearly specify which vulnerability it addresses in each response. As emphasized by prior studies (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024), explicitly prompting LLMs to articulate their reasoning has shown significant improvements in the accuracy and reliability of their outputs. Furthermore, clear explanations from the code author agent enable other agents (the moderator and review board) to better understand the justification or rationale behind the agent’s position. For the security researcher agent, this reasoning would be used for building their rebuttal response. This structured approach helps ensure that subsequent agents are equipped to offer well-informed counterarguments, summaries, or final evaluations.
You are the Code Author of the attached code.
The Security Researcher has presented a JSON array of alleged vulnerabilities. You must respond as if you are presenting your case to a group of decision-makers who will evaluate each claim. Your tone should be respectful, authoritative, and confident, as if you are defending the integrity of your work to a panel of experts.
For each identified vulnerability, produce a corresponding JSON object with the following fields:
vulnerability”: The same name/description from the Security Researcher’s entry. response-type”: ‘refutation’ if you believe this concern is unfounded, or ‘mitigation’ if you acknowledge it and propose a workable solution. reason”: A concise explanation of why the vulnerability is refuted or how you propose to mitigate it. ¡code¿
¡chat history¿
Figure 4. Code author agent prompt.
2.3. Moderator Agent (Judge Role)
In VulTrial, the moderator agent attempts to mirror the role of a judge in a courtroom scenario. Just as judges provide guidance to jurors and uphold fairness, within the context of vulnerability detection, this agent is tasked with objectively summarizing the arguments presented by both the security researcher agent (prosecutor) and the code author agent (defense attorney). It should focus only on the facts, avoiding any bias toward either side’s party. The primary purpose of this impartial summary is to assist the review board agent in reaching a decision, while also clarifying critical discussion points for subsequent rounds of argumentation by the security researcher and code author agents.
The prompt for this agent is highlighted in Fig. 5. In addition to accessing the code snippet in question, the moderator receives the security researcher’s identified vulnerabilities and the code author’s counterarguments (included in ¡code history¿). The moderator’s output consists of concise summaries of these two viewpoints, highlighting the key points from each. Fig. 2 presents an example of the moderator agent’s output. It summarizes the arguments from the security researcher (e.g., the number of potential vulnerabilities identified, their main reasons, and impacts) as well as from the code author, such as which vulnerabilities were acknowledged or refuted with the main reasoning.
You are the Moderator, and your role is to provide a neutral summary.
After reviewing both the Security Researcher’s identified vulnerabilities and the Code Author’s responses, provide a JSON object with two fields:
security researcher summary”: A concise summary of the vulnerabilities and reasoning presented by the Security Researcher. author summary”: A concise summary of the Code Author’s counterarguments or mitigation strategies. ¡code¿
¡chat history¿
Figure 5. Moderator agent prompt.
2.4. Review Board (Jury Role)
The review board agent in this VulTrial is similar to the role of jury in a courtroom scenario. In a traditional courtroom, the jury weighs all evidence and testimony presented during a trial, alongside any guidance from the judge, to determine whether a defendant is guilty or not. Similarly, in vulnerability detection, the review board agent is tasked to evaluate the code’s security status based on the evidence presented by both the security researcher and code author agents, as well as the moderator’s summary. Its purpose is to deliver a final verdict on whether the potential vulnerabilities identified by the security researcher are indeed valid. The prompt guiding the review board agent is shown in Figure 6. This agent receives (1) the summary from the moderator agent; (2) the list of potential vulnerabilities from the security researcher agent; (3) the code author agent’s counterarguments; (4) the original code snippet. The review board agent is tasked with generating several outputs.
First, it must extract the names of potential vulnerabilities from the list provided by the security researcher and proceed with a decision for each one. The review board agent needs to highlight the specific vulnerability being discussed as there may be multiple vulnerabilities under review. Second, the agent must determine whether each identified vulnerability is “valid,” “invalid,” or “partially valid.” This process is similar to a jury deciding whether a defendant is guilty or not guilty. Third, the agent must assess the severity of the vulnerability, similar to how a jury determines the extent of damages in a trial. However, instead of assigning an open-ended value, the review board needs to categorize severity as “low,” “medium,” “high.” If the vulnerability is deemed invalid, the severity is marked as “none.” Fourth, the agent must provide a recommended action based on the decision and severity level. For example, if a vulnerability is deemed valid and has a high severity, the recommended action should be “fix immediately.” Finally, the agent must generate an explanation for its decisions. Unlike a jury, which is not required to provide reasoning, we ask the review board agent to do so, as asking LLMs for reasoning can enhance performance (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024) and helps users to understand the decision.
We employ a threshold-based approach to classify a function as vulnerable, rather than a simple binary “vulnerable or not,” because it offers flexibility for users to tailor vulnerability detection to their tolerance for risk and false positives. By adjusting the threshold, users can decide how strict they want the classification to be. For instance, if a user prefers to catch even the smallest chance of vulnerability (and is comfortable with a higher rate of false positives), they can opt to flag any partially valid vulnerabilities regardless of severity. Conversely, if a user wants fewer false positives, they can raise the threshold to consider only higher-severity issues as triggers.
In our study, we flag a function as vulnerable if at least one vulnerability raised by the security researcher agent is deemed valid at high severity with recommended action fix immediately by the review board agent. This stricter threshold ensures that only the critical and well-substantiated vulnerabilities contribute to the classification, reducing false positives from lower-risk or ambiguous issues. We aim to reduce noise by excluding vulnerabilities marked as partially valid or those with lower severity levels. Fig. 2 also includes an example of the review board agent’s output. It classifies the first vulnerability (Out-of-Bounds Read) as “valid” with “high” severity. Meanwhile, the second vulnerability (Type Handling) is marked as “partially valid” with “medium” severity. Since at least one confirmed vulnerability is present, the final verdict for this code is vulnerable.
You are the Review Board. After reviewing the Moderator’s summary, Code Author’s and Security Researcher’s argument, and ¡code¿, produce a JSON array of verdicts for each vulnerability identified by the Security Researcher. Each object in the array should include: vulnerability”: The same name as given by the Security Researcher. decision”: One of ‘valid’, ‘invalid’, or ‘partially valid’. severity”: If valid or partially valid, assign a severity (‘low’, ‘medium’, ‘high’); if invalid, use ‘none’. recommended action”: Suggest what should be done next (e.g., ‘fix immediately’, ‘monitor’, ‘no action needed’). reason”: A brief explanation of why you reached this conclusion, considering both the Security Researcher’s and Code Author’s perspectives. ¡code¿
¡chat history¿
Figure 6. Review board agent prompt.
2.5. Agent Interaction
In VulTrial, four agents interact sequentially: the security researcher agent, the code author agent, the moderator agent, and finally the review board agent. During the first round, the security researcher agent (1) only has access to the original code snippet. In subsequent rounds, it also receives information from the code author and moderator. Similarly, the code author agent (2) initially sees only the security researcher agent’s findings, but in the next rounds it can also incorporate the moderator’s feedback. This setup is similar to a courtroom process, which can allow for multiple rounds of arguments and counterarguments between the prosecutor, defense attorney, and judge. The review board agent (4), mirroring a jury role, provides its verdict only after all discussions have concluded. It does not participate in the debate itself but bases its final decision on the evidence and summaries presented by the other three agents.
Allowing multiple discussion rounds may enhance the iterative refinement of analyses and arguments. It would also provide more chances for the security researcher agent and code author agent for rebuttal. Nevertheless, increased rounds entail higher operational costs. Therefore, it is essential to balance the trade-off between performance improvement and the cost. In this study, we initially limit the number of discussion rounds to one (see Section 4.1), and subsequently, in Section 4.3, we explore how varying the number of discussion rounds impacts the overall performance of VulTrial.
3. Experimental Setup
3.1. Dataset
In our study, we utilized the PrimeVul dataset, recently proposed by Ding et al. (Ding et al., 2025b), to evaluate LLMs using single-agent and multi-agents on vulnerability detection. Specifically, we used the PrimeVul pair data, which presents a challenging benchmark, as the vulnerable and benign code functions share at least 80% of their string content. This dataset ensures no data leakage or exact duplication between the training, testing, and validation sets. The latest version of PrimeVul Pair contains 435 vulnerable and 435 benign functions (Ding et al., 2025a).
3.2. Research Questions
This study aims to answer the following research questions (RQs):
RQ1: How does the performance of VulTrial compare to that of single-agent and multi-agent baselines on PrimeVul? Prior work (Ding et al., 2025b) demonstrated the poor performance of GPT-3.5 and GPT-4 for vulnerability detection on pairwise vulnerable and benign functions in PrimeVul. These findings highlight the limitations of single-agent prompting strategies in challenging detection scenarios. In this RQ, we investigate whether VulTrial, a multi-agent LLM-based framework (Section 2), can outperform both single-agent methods (Ding et al., 2025b) and existing multi-agent approaches such as GPTLens (Hu et al., 2023) on PrimeVul.
RQ2: How does instruction tuning on role-specific agents influence the performance of automated vulnerability detection? VulTrial consists of four agents, each with a specific role. Instruction tuning can be applied to each agent to strengthen its performance in its designated task. In this RQ, we investigate whether this role-specific instruction tuning can enhance the overall performance of VulTrial.
RQ3: Does allowing multiple rounds of debate among role-specific agents in VulTrial lead to improvement compared to a single-pass analysis? In previous RQs, we focus on conducting a single round of discussions among agents to reach a final verdict. However, allowing multiple rounds of discussion could give agents the opportunity to refine their arguments to reach a more accurate decision. In this RQ, we examine whether enabling multiple rounds of discussion can enhance VulTrial’s performance.
RQ4: How does each agent contribute to the overall performance of VulTrial VulTrial features four distinct roles, security researcher, code author, moderator, and review board, which is designed to follow a courtroom-style scenario. While the results from previous RQs highlight VulTrial’s effectiveness, it is not known whether each agent is essential to the final prediction and how it impacts the performance. In this RQ, we do the ablation study that systematically omits one agent at a time to help us verify each agent’s contribution and evaluate whether the loss of a role reduces the performance.
Important note: The *ground-truth label* for this code is: “¡vulnerable/benign¿ vulnerable” means this code contains at least one vulnerability. benign” means there are no vulnerabilities in this code. Use the label to ensure your arguments match the fact: Even if you are “defending” the code, if the label is “vulnerable,” you should not fully deny actual issues. If the label is “benign,” you can confidently refute all vulnerability claims. Here is the CVE ID and description that is related with the code:
¡CVE ID¿”: ¡CVE Description¿
Figure 7. Example of additional information provided to the code author agent for constructing the instruction tuning dataset.
3.3. Baseline and Models
Baselines. For the single-agent baseline, we utilized the chain of thought (CoT) (Wei et al., 2022) prompting technique, which achieved the highest results in the PrimeVul (Ding et al., 2025b). CoT prompting encourages LLMs to reason through problems step by step to arrive at an answer. For the multi-agent baseline, we utilized an approach from a recent previous study, GPTLens (Hu et al., 2023). GPTLens operates with two types of agents: auditor agent, which identifies potential vulnerabilities, and a critic agent, which evaluates and scores the identified vulnerabilities. The final decision is based on the critic’s score; if the score is less than 5, it is classified as benign; otherwise, it is considered vulnerable. There are several key differences between GPTLens and VulTrial. VulTrial follows a courtroom-inspired setting with four role-specific agents, whereas GPTLens relies on a two-agent structure. Unlike GPTLens, where the critic agent determines the final decision, VulTrial used a different agent (i.e., review board agent) for this task, designed to be neutral toward the security researcher and code author agents. Recent studies (Gao et al., 2024; Hu et al., 2024) suggest that employing domain-specific multi-agent systems that reflect a more realistic scenario leads to better decision making, further supporting our approach in VulTrial.
Models. We employed multiple models in this study. First, we used several well-established OpenAI models known for their strong performance, as confirmed by prior research (Widyasari et al., 2024; Hu et al., 2023; Zhou et al., 2024). Specifically, we utilized GPT-3.5 and GPT-4o.
GPT-3.5 (OpenAI, 2023a): GPT-3.5 is one of the first widely popular OpenAI LLM models. Following PrimeVul (Ding et al., 2025b), we also evaluate the performance of this model in our study. In our evaluation, we utilized gpt-3.5-turbo-0125, the latest iteration of the GPT 3.5 model with a context window of 16k tokens.
GPT-4o (OpenAI, 2024b): We employed GPT-4o, one of OpenAI’s flagship large language models recommended as a replacement for GPT-4 (OpenAI, 2023b). We chose GPT-4o over GPT-4 (used in PrimeVul’s evaluation) because OpenAI indicates that GPT-4o offers performance comparable to GPT-4 while significantly reducing computational costs (OpenAI, 2024b). Specifically, we used gpt-4o-2024-08-06 allows up to 128k tokens in its context window and 16k output tokens.
3.4. Evaluation Metrics
Following the approach in PrimeVul (Ding et al., 2025b), we adopt a pair-wise evaluation strategy. It is designed to assess the model’s prediction on pairs of vulnerable and benign code, rather than the single codes. This method highlights the model’s capacity to differentiate subtle variations that make one function vulnerable and the other benign. The four key metrics are defined as follows:
  • Pair-wise Correct Prediction (P-C): The model correctly predicted both codes in the pair.
  • Pair-wise Vulnerable Prediction (P-V): The model incorrectly classifies both code as vulnerable.
  • Pair-wise Benign Prediction (P-B): The model incorrectly classifies both code as benign.
  • Pair-wise Reversed Prediction (P-R): The model assigns the opposite label to each code in the pair, predicting “benign” for a vulnerable function and “vulnerable” for a benign one.
We use P-C as our primary evaluation metric. A high P-C indicates accurate vulnerability detection that avoids false positives and negatives, making it a more robust indicator of overall performance.
3.5. Methods
3.5.1. Answering RQ1:
We compare VulTrial with both single-agent (Ding et al.’s CoT(Ding et al., 2025b)) and multi-agent (GPTLens) baselines, using two different LLMs, i.e., GPT-3.5 and GPT-4o. Section 3.3 provides further details on our baseline methods and model choices. We evaluate each approach using several metrics, with the Pair-wise Correct (PC) rate as our primary evaluation metric. A higher PC rate indicates better performance.
3.5.2. Answering RQ2:
For this RQ, we conducted instruction tuning using samples drawn from GPT-4o’s validation set. After running GPT-4o on the validation set, we took its correct outputs to create an instruction tuning dataset for a GPT-4o model. We selected GPT-4o because, as shown in RQ1, VulTrial using GPT-4o outperforms both single-agent and multi-agent baselines and achieves the best overall results, making it an ideal candidate for evaluating the impact of instruction tuning. Our goal is to determine whether instruction tuning can further enhance its performance. We chose GPT-4o for generating the instruction data to ensure consistency in style and reasoning patterns. By using the same models for generating data, we reduce the risk of mismatches that can occur when instruction tuning with outputs from a completely different model.
For instruction tuning data, we use 50 code pairs (vulnerable and benign) with the fewest tokens in the validation set to help manage the tuning cost. In addition, OpenAI has highlighted that GPT-4o can achieve strong results with “as few as a few dozen examples” (OpenAI, 2024a). When creating our instruction tuning dataset, we initially included the true label in the prompt (e.g., adding vulnerability-related information for the code author agent, see Fig. 7), but then removed direct label disclosures in the final dataset. As a result, each prompt retained its original role-specific structure and was paired with the correct assistant output.
We utilized following scenarios for the evaluation: (1) no instruction tuning, (2) instruction tuning only the security researcher agent, (3) instruction tuning only the code author agent, (4) instruction tuning only the moderator agent, (5) instruction tuning only the review board agent, (6) instruction tuning both security researcher and code author agents, (7) instruction tuning the security researcher, code author, and moderator agents, and (8) instruction tuning all agents. To systematically assess the impact of instruction tuning, we first applied it to individual agents to understand their specific contributions. Then, we incrementally introduced instruction tuning in the order of agent interaction, allowing us to analyze how tuning at different stages influences the overall performance of VulTrial. This stepwise approach provides insights into the cumulative effects of instruction tuning across multi-agents.
3.5.3. Answering RQ3:
We experimented with three values of $k$ (i.e., the number of interaction rounds): 1, 2, and 3 inVulTrial. We selected these values to systematically investigate the impact of multiple interaction rounds while maintaining computational feasibility. For the base model, we used GPT-4o with instruction tuned agent, as it yielded the strongest overall performance.
3.5.4. Answering RQ4:
In this research question (RQ), we conduct an ablation study by systematically removing each agent to assess its necessity. Since the security researcher agent is fundamental for identifying potential vulnerabilities, we retain it in all scenarios. Additionally, since the moderator primarily provides summaries rather than making explicit decisions on vulnerabilities, we do not include a scenario where the moderator is responsible for delivering the final verdict.
In total, we evaluate four scenarios. First, the Full VulTrial (baseline) includes all four agents. Second, we exclude the moderator and use only the security researcher, code author, and review board. Third, we exclude the code author. Finally, we consider the effect of removing two agents at once. In this fourth scenario, we explore (a) only the security researcher and review board (i.e., no code author, no moderator), and (b) only the security researcher and code author (i.e., no moderator, no review board). For scenario 4(b), as the final decision is made by the code author agent, we classify the code as vulnerable if the code author mitigates at least one of the potential vulnerabilities.
By comparing performance across these scenarios, we can identify which agents play the most critical roles. For example, does a moderator assist the review board in reaching a final decision, and how essential is a different agent for delivering verdicts? This ablation would highlight the necessity of each agent in VulTrial.
Table 1. Results of various LLMs on PrimeVul paired functions. An upward-pointing arrow ($\uparrow$) indicates that a higher score is better, while a downward-pointing arrow ($\downarrow$) indicates that a lower score is better.
Model Agents Method P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
Ding et al.’s CoT (Ding et al., 2025b) 6.21 4.79 83.51 5.5
Single-Agent Finetuned 1.24 5.32 90.96 2.48
GPTLens (Hu et al., 2023) 4.60 89.20 0.69 5.51
GPT-3.5 Multi-Agent VulTrial 15.63 9.20 60.92 14.25
Single-Agent Ding et al.’s CoT (Ding et al., 2025b) 9.20 9.89 74.26 6.67
GPTLens (Hu et al., 2023) 10.11 55.40 28.05 6.44
GPT-4o Multi-Agent VulTrial 18.62 41.15 28.74 11.49
  • $\dagger$ Results taken from previous study by Ding et al. (Ding et al., 2025b).
3.6. Implementation Details
We built VulTrial on top of AgentVerse (Chen et al., 2024), a recently developed framework that simplifies the deployment of multiple LLM-based agents. For single-agent baselines, we used the prompt in PrimeVul (Ding et al., 2025b), while for GPTLens (Hu et al., 2023), we employed the provided replication package (Hu et al., 2022). Because the original GPTLens study focused on smart contracts, we adapted its prompt by replacing references to “smart contract” with “code” to generalize the use case. We used the best settings reported in their study, including two auditor agents, a top-$k$ value of 3, and the criteria for vulnerable with a final score of at least 5 (out of 9). All GPT-based models were accessed through the OpenAI API. For instruction tuning GPT-3.5 and GPT-4o, we utilized OpenAI’s official fine-tuning tools (OpenAI, 2024a). To make the results more deterministic, we set the temperature to 0 for both our single-agent baselines and VulTrial. For GPTLens (Hu et al., 2023), we followed the same setting used in its evaluation (0.7 for the auditor agent and 0 for the critic agent).
4. Results
4.1. RQ1: VulTrial vs. Single-Agent and Multi-Agent Baselines
In this RQ, we evaluated VulTrial against single-agent (using Ding et al. CoT prompts (Ding et al., 2025b)) and multi-agent (GPTLens (Hu et al., 2023)) baselines across several LLMs. We utilized GPT-3.5 and GPT-4o for the base model in VulTrial, single-agent, and multi-agent baselines (see Section 3). Table 1 presents the results, with P-C indicating pairs classified correctly for both functions (higher is better).
GPT-3.5.
Among all models, GPT-3.5 saw the largest boost when applying VulTrial. Its P-C score increased by 151.69%, over the single-agent baseline. Likewise, P-B dropped from 83.51 to 60.92, which highlights the number of both codes in the pair that were predicted as benign are less in VulTrial. Furthermore, VulTrial also improves upon the performance of GPTLens by 239.78% on P-C. In most cases, GPTLens predicts functions as vulnerable, as shown P-V score (i.e., pairs incorrectly labeled as vulnerable) of 89.2, which is much higher compared to VulTrial (9.2).
GPTLens is configured to output the top-3 potential vulnerabilities per function, following the best setup from the original study (Hu et al., 2023). Since auditor agent runs twice, this results in up to six potential vulnerabilities per function. Indeed, most auditor outputs list six vulnerabilities, with an average of 5.6 and a median of 6. Meanwhile, VulTrial produce an average of 1.74 vulnerabilities per function, with a median of 2. In other words, VulTrial produces fewer vulnerabilities compared to GPTLens, but it has more precise vulnerability detection, shown by a higher P-C.
GPT-4o.
Using GPT-4o as the base model, VulTrial also achieves the best performance. Specifically, it outperforms the single-agent baseline by 102.39% and the multi-agent baseline by 84.17% in the P-C metric. Across all methods (VulTrial, single-agent, and multi-agent), performance under GPT-4o is higher than with GPT-3.5. This aligns with OpenAI’s claim that GPT-4o offers higher performance than GPT-3.5, as well as with previous benchmarking studies that compare these models (OpenAI, 2024b; Wang et al., 2024; Günay-polatkan and Sığırlı, 2024; Kipp, 2024).
For the single-agent baseline with GPT-4o, similar to its GPT-3.5 performance, it exhibited a high P-B score (pairs incorrectly classified as benign), suggesting a tendency to overlook vulnerabilities. However, in GPT-4o the P-B score dropped to 74.26, compared to GPT-3.5 (83.51). In contrast, VulTrial significantly reduced P-B to 28.74, a lower value than both the single-agent GPT-4o and its own performance under GPT-3.5 (60.92). This demonstrates VulTrial’s improved ability to correctly predict benign code with GPT-4o. Examining P-R (pairs where the predicted classification is reversed from the ground truth), VulTrial with GPT-4o also showed an improvement, reducing P-R by 24.02% compared to its performance with GPT-3.5.
In the GPTLens experiment using GPT-4o, the first agent highlights a similar number of potential vulnerabilities as it did under GPT-3.5, with a median of 6 and a mean of 5.76. Meanwhile, VulTrial with GPT-4o reports fewer vulnerabilities (median 4, mean 4.11) compared to GPTLens. Despite highlighting fewer vulnerabilities, VulTrial achieved higher P-C and lower P-B scores than GPTLens, indicating overall better performance. This suggests that the additional agents in VulTrial, can effectively distinguish between justified and unjustified vulnerability claims, leading to more accurate final predictions.
In Fig. 2, we show an example conversation from VulTrial. The code snippet under discussion is flagged for failing to validate negative indices, which allow an attacker to read arbitrary data by inputting out-of-range values in indices. The security researcher agent identifies two vulnerabilities: (1) Out-of-Bounds Read (CWE-125), which is valid, and (2) Type Handling, which is misleading since the existing code already handles unsupported types safely. After reviewing these findings, the code author agent agrees with the Out-of-Bounds Read assessment but refutes the Type Handling concern, which shows the agent can distinguish confirmed risks from exaggerated claims. The moderator agent then summarizes key points from both agents, which aim to help review board agent in making the decision. Based on these information, the review board agent concludes that the Out-of-Bounds Read is a genuine vulnerability. Meanwhile, for the Type Handling issue, the review board agrees with the code author’s argument that the existing implementation is safe. However, they worried about potential future changes that could introduce vulnerabilities as highlighted by the security researcher agent. As a result, the board classifies this issue as partially valid with a recommendation to monitor.
Answer to RQ1 Using GPT-3.5 as base LLMs, VulTrial improves P-C of single-agent and multi-agent baseline by 151.69% and 239.78%, respectively. Meanwhile, using GPT-4o as the base LLMs, VulTrial improve the baselines by 102.39% and 84.17%.
4.2. RQ2: Instruction Tuning on Role-Specific Agents
From the RQ1’s results, we observed that VulTrial using GPT-4o achieve the best performance. In this RQ, we investigate whether instruction tuning on role-specific agents can further improve VulTrial’s overall performance. The results of this instruction tuning phase are shown in Table 2.
Table 2. Results of VulTrial using GPT-4o with instruction tuned agents. SR represents the security researcher agent, CA the code author agent, M the moderator agent, and RB the review board agent.
Instruction Tuning P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
None 18.62 41.15 28.74 11.49
Security researcher (SR) 12.41 35.17 42.99 9.43
Code author (CA) 11.95 8.97 72.18 6.90
Moderator (M) 22.07 35.40 30.80 11.72
Review board (RB) 16.55 30.57 42.07 10.80
SR + CA 7.82 9.20 76.09 6.90
SR + CA + M 7.36 8.51 77.70 6.44
SR + CA + MR + RB 10.34 10.57 70.11 8.97
GPT-4o Instruction Tuning.
All the single instruction tuned variants of GPT-4o in VulTrial, covering the security researcher agent, code author agent, moderator agent, and review board agent, outperformed both the single-agent and multi-agent baselines. Among these, tuning the moderator agent delivered the most substantial performance gain. Specifically, the moderator-tuned version increases the performance of the non-tuned version in VulTrial by 18.53% (P-C). Compared to the single-agent baseline, this instruction tuned version of VulTrial increases P-C by 139.89%. While for the multi-agent baseline, the improvement of the instruction tuned version of VulTrial increases P-C by 118.30%.
Effect of Agent-Specific Tuning.
Upon further analysis of the instruction tuned moderator agent, we observed that its summaries are more concise compared to the non-tuned version. Specifically, the tuned version of the moderator agent provides a shorter summary than the non-tuned version, reducing the summary length for security researchers by 16.03% and for code authors by 16.58%. In cases where the tuned VulTrial correctly classified code as benign but the non-tuned version did not, the security researcher’s summary was 21.31% shorter in the tuned version. Meanwhile, when the tuned VulTrial correctly classified code as vulnerable but the non-tuned version did not, the code author’s summary was 19.03% shorter. These observations suggest that the instruction tuned moderator agent can effectively distill key points from both the security researcher and code author, which help the review board agent to make more accurate final verdicts.
We next examined the impact of instruction tuning on the security researcher and code author agents. Although both configurations still surpassed single-agent and prior multi-agent baselines, their performance fell below that of the non-tuned VulTrial. After tuning, we found that the security researcher agent proposed fewer potential vulnerabilities, with an average of 2.58 per code instance compared to 4.11. To generate the tuning data, GPT-4o is provided with ground truth information, such as CVE details. Since it already knows the ground truth vulnerabilities, the number of potentially vulnerable instances identified by security researcher agent is smaller. On average, security researcher identify 1.25 vulnerabilities in the training data, which is lower compared to the non-tuned version. This low number of vulnerabilities in the tuning data leads to a lower number of vulnerabilities identified in the tuned version of security researcher agent.
Similarly, instruction tuning the code author agent altered its mitigation-to-refutation ratio. While the mitigation to refutation ratio of non-tuned version was 2:1, the tuned version exhibited a 1:4 ratio. In other words, the tuned code author refuted vulnerabilities far more frequently. From its tuning data, it learned to doubt potential vulnerabilities highlighted by security researcher. Although this approach can restrict over-diagnosing vulnerabilities, it also risks undervaluing genuine security concerns, as highlighted by an increase in P-B (cases where both snippets are incorrectly labeled benign). Despite outperforming single-agent baseline, this shows the delicate balance that has to be achieved by instruction tuning; While instruction tuning can sharpen an individual agent’s skills, it may introduce biases or extreme behaviors that undermine collective performance.
Combining Agent-Specific Tuning.
We also explored configurations with multiple tuned agents, such as the security researcher and code author, are used together in VulTrial. These combinations did not surpass the performance achieved by tuning only the moderator agent. For instance, when all agents were instruction tuned, the average number of identified vulnerabilities dropped from 4.11 to 2.60, and the refutation-to-mitigation ratio shifted to 1:4. This imbalance corresponded to a higher P-B rate of 70.11.
Answer to RQ2 Instruction tuning for moderator agent using a small dataset (50 paired samples) yields the most significant improvements. It enhances performance by 18.53% compared to the non-tuned version and increases P-C by 139.89% and 118.30% over the single-agent and multi-agent baselines, respectively.
4.3. RQ3: Effect of Multiple Rounds of Discussion
In the previous RQs, VulTrial used a single round of interaction among the agents. In this RQ, we investigate whether additional discussion rounds could enhance performance. The results of these multiple rounds experiments are shown in Table 3. The best result is achieved by using a single round interaction.
Table 3. Results of VulTrial using GPT-4o across different numbers of discussion rounds.
Model #Round P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
1 22.07 35.40 30.80 11.72
2 20.00 39.77 28.05 12.18
GPT-4o 3 18.62 44.60 25.29 11.49
  • $\ddagger$ Instruction tuning for moderator agent.
For GPT-4o, increasing the number of rounds actually led to a decrease in overall performance. Specifically, the proportion of pairs incorrectly categorized as vulnerable (P-V) increases with more rounds. From further analysis, we found that the security researcher agent suggested fewer potential vulnerabilities in later rounds, with an average of 4.22 per code in Round 1, decreasing to 3.55 in Round 2 and 3.43 in Round 3. This decline is also driven by the code author agent’s refutations. In 30.88% out of the total prediction, the security researcher removed a previously identified vulnerability after the code author disputed it. Meanwhile, the code author agent became progressively more inclined to acknowledge vulnerabilities, with its mitigation-to-refutation ratio shifting from 1.73:1 in Round 1, to 2.5:1 in Round 2, and finally 3:1 in Round 3. Although the review board agent did not blindly accept arguments from either side (for instance, during the three-round setup, in 100 out of 868 code instances, review board agent overruled the code author’s refutation and decided the vulnerability to be valid), the overall imbalance introduced by these interactions still raised P-V and lower P-C rates.
Answer to RQ3 Using GPT-4o in VulTrial outperforms both single-agent and multi-agent baselines across different rounds of agent interaction (i.e., 1, 2, and 3). However, performance declines when extending beyond a single discussion round.
Table 4. Results of ablation study on agent in VulTrial using GPT-4o with moderator agent tuned.
Agents P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
Base 22.07 35.40 30.80 11.72
No moderator 7.36 8.51 77.70 6.44
No code author 14.48 63.68 11.95 9.89
No moderator+code author 14.48 61.84 14.48 9.20
No moderator+review boards 2.07 93.56 1.15 3.22
4.4. RQ4: Ablation Study
In this RQ, we investigate the necessity of each agent in VulTrial. Table 4 presents the ablation results under different agent-removal scenarios. Overall, we find that all four agents are essential for achieving the best performance.
Removing the moderator agent from VulTrial reduces P-C from 22.07 to 7.36. This result highlights the moderator’s critical role in summarizing arguments to help review board agent in making the decision. This finding is also consistent with the results of RQ2 (see Section 4.2), where instruction tuning the moderator yielded the largest performance gain. Removing both the moderator and review board agent causes P-C to further drop 2.07, while P-V (pairs both labeled vulnerable) increase to 93.56. Highlighting that most of the codes are predicted as vulnerable by the code author. This imbalance occurs because, without an additional agent review board, the final judgment defaults heavily toward labeling codes as vulnerable based on partial agreements from the code author agent. Thus, the review board is essential for ensuring balanced final verdicts.
Removing the code author agent from VulTrial reduces P-C to 14.48 and increases P-V to 63.68. Without the code author’s counterarguments, the review board agent has a limited perspective and frequently accepts vulnerabilities identified by the security researcher agent. Removing both the moderator and code author produces similar results (P-C: 14.48, P-V: 61.84). This underscores the crucial role of the code author in ensuring balanced judgments while also demonstrating that the moderator agent becomes redundant without opposing arguments to mediate.
Answer to RQ4 Our findings confirm that every agent in VulTrial is necessary to achieve the best performance.
5. Discussion
Cost vs. Performance Trade-offs
By assigning specialized roles to agents, mirroring real-world scenarios, VulTrial can facilitate deeper discussions by incorporating diverse rationales from different parties, even when using small-scale LLMs. However, this multi-agent approach inherently incurs additional costs due to the increased number of input and output tokens. Interestingly, while GPT-4o is generally more expensive and performs better than GPT-3.5 as a single-agent, with respective P-C 9.2 and 6.21, using a multi-agent approach (VulTrial) with GPT-3.5 outperforms GPT-4o’s single-agent performance by 69.89%. This approach also reduces costs, requiring only \$7.46 compared to GPT-4o’s single-agent cost of \$8.10 for the whole testing data. Conversely, applying VulTrial to more expensive models, such as GPT-4o, significantly increases costs (approximately \$23.9=32-8.1). Therefore, careful consideration is important when selecting models to optimize both cost efficiency and detection performance.
Multi-Agents Performance
In both experiments using GPT-3.5 and GPT-4o as base models for GPTLens (see Section 4.1), we observed that a large portion of the codes were predicted as vulnerable, as shown by high P-V scores (incorrectly predicted as vulnerable) of 89.2 and 41.15, respectively. This behavior may be due to the critic agent (final agent) assigning high scores on potential vulnerabilities flagged by the auditor agent (first agent). To get the final prediction, GPTLens also utilizes a threshold to determine whether a function is vulnerable. The final score, ranging from 0 to 9, is computed as $0.5\times\text{correctness}+0.25\times\text{severity}+0.25\times\text{profitability}$, with correctness, severity, and profitability taken from the critic agent output.
The final scores below 5 are considered benign. When GPTLens uses GPT-3.5, the median final score is 7.25 with mean 7.08, indicating that most code is labeled as vulnerable. Similarly, when GPTLens uses GPT-4o, the median final score is 6.75 with mean 6.67, which means the majority of the code prediction also exceeds the benign threshold.
Meanwhile, under GPT-4o, VulTrial demonstrated a more balanced results, as indicated by a mitigation-to-refutation ratio of 2:1 in the code author agent’s responses. Furthermore, we found that in 101 benign and 253 vulnerable codes, the code author agent refuted at least one potential vulnerability initially flagged by the security researcher (i.e., disagreement between code author and security researcher), yet the final label remained correct. These cases, accounting for 40.78% of the total codes, demonstrate VulTrial’s effectiveness in resolving disagreements.
Enhancing Role-Specific Agent
Tuning the moderator agent in VulTrial using GPT-4o enhances single-agent performance by 139.89% and improves the non-tuned version of VulTrial by 18.52%. This underscores the importance of instruction tuning for role-specific agents. While instruction tuning on all four agents in VulTrial would cost four times more than tuning a single-agent. In this study, we used only a sample of 50 pairs for instruction tuning, which kept the cost at approximately \$17 for each agent. This is significantly lower than using the instruction tuning setting done in PrimeVul (Ding et al., 2025b), which would cost approximately more than \$500 for GPT-4o (¿15,000 samples).
6. Threats to Validity
External Validity. An external threat to our study is the generalizability of our findings. In this study, we focused the evaluation using PrimeVul pair data, which is a challenging recently-released vulnerability detection task on C/C++ programming language. We did not use other datasets as they may not be free from bias that have been uncovered by Ding et al. (Ding et al., 2025b). In the future, we plan to investigate VulTrial performance using other clean datasets.
Construct Validity. Threats to construct validity are from the choice of metrics in our study. In this study, we utilize P-C, P-B, P-V, P-R (Section 3.4). These metrics are designed for a pair-wise evaluation introduced in PrimeVul (Ding et al., 2025b). In a textually similar pair of functions, the metrics assesses an approach’s ability to correctly identify both the presence and absence of vulnerabilities. Consequently, we believe that the threat to construct validity is minimal.
Internal Validity. Threats to internal validity in our study refer to possible errors in our experiment. To mitigate the risk, we have carefully checked the experiments that we run in our study. The code, prompt, and results are also included in our replication package: https://figshare.com/s/1514bc9a7aa64b46d94e. Furthermore, to reduce the randomness effect of the experiment, we set the temperature to 0 to make the results more deterministic. We also did the preliminary evaluation where we ran the VulTrial on testing data three times and found that the standard deviation is low (0.7).
7. Related Work
Single-Agent Vulnerability Detection. A broad range of studies have explored prompting and CoT for vulnerability detection (Tamberg and Bahsi, 2024; Ullah et al., 2023; Zhou et al., 2024; Nong et al., 2024; Zhang et al., 2024; Ding et al., 2025b; Steenhoek et al., 2024). Tamberg and Bahsi (Tamberg and Bahsi, 2024) benchmarked LLMs using advanced prompting techniques, including CoT, tree-of-thought, and self-consistency, finding that CoT with GPT-4 achieved the best performance and outperformed static analysis tools. Similarly, Zhou et al. (Zhou et al., 2024) demonstrated that GPT-4 with prompting outperforms fine-tuned CodeBERT. Ullah et al. (Ullah et al., 2023) evaluated commercial LLMs across various prompting strategies, including CoT, and found promising detection performance. However, they noted often unfaithful reasoning in LLMs leads to incorrect classifications. Recent studies have further highlighted the limitations of CoT (Nong et al., 2024; Zhang et al., 2024; Steenhoek et al., 2024). Notably, PrimeVul’s seminal work found that GPT-3.5 and GPT-4 perform poorly on PrimeVul’s pair test set, reinforcing concerns about CoT’s reliability in vulnerability detection. In this work, we use CoT as a baseline and show that our approach, VulTrial, significantly outperforms it across various LLMs on PrimeVul.
Multi-Agent Vulnerability Detection. Multi-agent frameworks can enhance LLM-based vulnerability detection by enabling collaboration and specialization among agents (He et al., 2024). For example, EvalSVA (Wen et al., 2024) introduces a multi-agent framework where agents act as security experts that improves over single-agent LLMs. Similarly, MuCoLD (Mao et al., 2024) adopts a multi-agent setup where agents assume roles such as tester and developer, which collaboratively analyze code. In contrast, VulTrial employs a structured multi-agent design inspired by a courtroom, where agents take on distinct roles such as prosecutor, defender, and judge. This courtroom analogy encourages adversarial debate and balanced reasoning that leads to robust decision-making. We excluded these recent works as baselines because the authors did not provide a replication package for reusing their agents, and the prompt information in the paper was insufficient to replicate the study.
Other multi-agent-based approaches have been explored for smart contract auditing (Wei et al., 2024; Ma et al., 2024; Hu et al., 2023). LLM-SmartAudit (Wei et al., 2024) uses collaborative decision-making with role-play and reasoning to detect vulnerabilities in smart contracts. Similarly, iAudit (Ma et al., 2024) assigns specific roles, such as detector, reasoner, and critic, to guide the analysis of Solidity smart contracts. GPTLens (Hu et al., 2023) employs multiple LLM-based agents acting as independent prosecutors to identify vulnerabilities, with a critic agent reviewing and ranking the findings. Our work differs from these studies in two key ways. First, we focus on vulnerability detection in C and C++ on PrimeVul, a benchmark comprising challenging real-world examples. Second, VulTrial adopts a structured, courtroom-inspired framework specifically tailored for general-purpose languages beyond smart contracts. As these studies focus on smart contracts, their scope differ from our work for general-purpose programming languages. Nonetheless, we found that GPTLens can be adapted to general-purpose vulnerability detection and included it as a baseline in our experiments (Section 3.3). Our results demonstrate that VulTrial substantially outperforms GPTLens and single-agent baselines on PrimeVul.
8. Conclusion and Future Direction
In this study, we proposed VulTrial, a novel multi-agent LLM-based approach for vulnerability detection that applies a courtroom-style paradigm. VulTrial comprises of four specific-role agents with distinct roles: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). Each agent is essential to get the best performance (Section 4.4). We evaluated VulTrial against single-agent (Ding et al.’s CoT) (Ding et al., 2025b) and multi-agent, GPTLens (Hu et al., 2023), vulnerability detection baselines. Our results show that VulTrial outperforms both the single-agent and multi-agent baselines. In experiments using GPT-3.5, VulTrial enhances the performance of the single-agent baseline by 151.69%, and outperforms GPTLens by 239.78%. Additionally, instruction tuning with just 50 sample pairs yields further performance improvement. When utilizing GPT-4o with instruction tuning in moderator agent, VulTrial improves single-agent and multi-agent baselines by 139.89% and 118.3%. Furthermore, the number of discussion rounds among the first three agents in VulTrial does not improve the overall performance.
For future direction, we plan to apply VulTrial to different domains to evaluate its generalizability further. We also plan to test additional LLMs such as DeepSeek-R1 (DeepSeek-AI et al., 2025) and LLaMA3 (Grattafiori et al., 2024) and consider each model’s strengths and weaknesses to choose the one that best fits for each role-specific task.

### v2

@@
 Detecting vulnerabilities in source code remains a critical yet challenging task, especially when benign and vulnerable functions share significant similarities.
-In this work, we introduce VulTrial, a courtroom-inspired multi-agent framework designed to enhance automated vulnerability detection.
+In this work, we introduce VulTrial, a courtroom-inspired multi-agent framework designed to identify vulnerable code and to provide explanations.
 It employs four role-specific agents, which are security researcher, code author, moderator, and review board.
-Through extensive experiments using GPT-3.5 and GPT-4o we demonstrate that VulTrial outperforms single-agent and multi-agent baselines.
-Using GPT-4o, VulTrial improves the performance by 102.39% and 84.17% over its respective baseline.
-Additionally, we show that role-specific instruction tuning in multi-agent with small data (50 pair samples) improves the performance of VulTrial further by 139.89% and 118.30%.
-Furthermore, we analyze the impact of increasing the number of agent interactions on VulTrial’s overall performance.
-While multi-agent setups inherently incur higher costs due to increased token usage, our findings reveal that applying VulTrial to a cost-effective model like GPT-3.5 can improve its performance by 69.89% compared to GPT-4o in a single-agent setting, at a lower overall cost.
-copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
+Using GPT-4o as the base LLM, VulTrial almost doubles the efficacy of prior best-performing baselines.
+Additionally, we show that role-specific instruction tuning with small quantities of data significantly further boosts VulTrial’s efficacy.
+Our extensive experiments demonstrate the efficacy of VulTrial across different LLMs, including an open-source, in-house-deployable model (LLaMA-3.1-8B), as well as the high quality of its generated explanations and its ability to uncover multiple confirmed zero-day vulnerabilities in the wild.
+vulnerability detection, large language models, multi-agent
+journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773256isbn: 979-8-4007-2025-3/2026/04
 1. Introduction
-Vulnerability detection in software code is a critical task to ensure system security and prevent potential exploits (Chen et al., 2023).
-Manual vulnerability detection, however, can be time-consuming and error-prone, especially given the scale and complexity of modern software systems.
-To automate this process, researchers have explored various machine learning and deep learning approaches (Zou et al., 2021a; Russell et al., 2018; Chernis and Verma, 2018; Risse and Böhme, 2024; Zou et al., 2021b; Chakraborty et al., 2021).
-Recently, with substantial advancements in large language models (LLMs) and their impressive performance on natural language tasks (Wang et al., 2023; Abbasiantaeb et al., 2024; Zhuang et al., 2023; Neumann et al., 2024), several studies have attempted to utilize LLMs for automated vulnerability detection tasks (Tamberg and Bahsi, 2024; Ullah et al., 2023; Zhou et al., 2024; Nong et al., 2024; Zhang et al., 2024; Widyasari et al., 2024), which shows promising outcomes.
+Vulnerability detection in software code is a critical task to ensure system security and prevent potential exploits (chen2023diversevul).
+Manual vulnerability detection can be time-consuming and error-prone, especially given the scale and complexity of modern software systems.
+To automate this, researchers have explored various machine learning and deep learning approaches (li2018vuldeepecker; russell2018automated; chernis2018machine; risse2024uncovering; zou2021interpreting; chakraborty2021deep).
+Recently, with substantial advancements in large language models (LLMs) and their impressive performance on natural language tasks (wang2023can; abbasiantaeb2024let; zhuang2023toolqa; neumann2024llm), several studies have found promising outcomes using LLMs for automated vulnerability detection (tamberg2024harnessing; ullah2023llms; zhou2024large_emerging; nong2024chain; zhang2024prompt; widyasari2024beyond)
 Despite these advances, recent research by Ding et al.
-(Ding et al., 2025b) highlighted significant limitations when evaluating current LLM-based vulnerability detection approaches.
+(ding2024vulnerability) highlighted significant limitations when evaluating current LLM-based vulnerability detection approaches.
 Their study proposed a challenging new benchmark called PrimeVul, where the functional differences between benign and vulnerable code pairs are subtle, less than 20%.
-These cases where the code has a high degree of similarity are critical to evaluate because in real-world scenarios, vulnerabilities often stem from subtle, non-obvious code changes (Ni et al., 2023; Tan et al., 2024).
+These cases where the code has a high degree of similarity are critical to evaluate because in real-world scenarios, vulnerabilities often stem from subtle, non-obvious code changes (ni2023distinguishing; tan2024similar).
 On this dataset, LLM-based approaches exhibited low performance.
-The best performance comes from using GPT-4 (OpenAI, 2023b), which correctly predicts both codes in a pair (vulnerable and benign), referred to as Pair-wise Correct Prediction (P-C), 12.4% of cases.
+The best performance comes from using GPT-4 (openai2023gpt4), which achieves a Pair-wise Correct Prediction (P-C) score of 40 (out of a possible 435).
+P-C measures how many vulnerable-benign code pairs a method classifies entirely correctly, directly reflecting its ability to detect the subtle code changes that introduce a vulnerability.
 This LLM-based approach used in PrimeVul was implemented under the conventional single-agent system, wherein one model attempts to identify and explain vulnerabilities on its own.
-Single-agent systems are inherently constrained by their reliance on a single perspective when analyzing vulnerabilities, which can lead to issues such as hallucination and overconfidence in incorrect reasoning (Zhang et al., 2023; Shi et al., 2025; Huang et al., 2025; Liang et al., 2023; Du et al., 2023).
-Moreover, these single-agent systems may fail to critically verify their own output, which is crucial when identifying subtle vulnerabilities (Liang et al., 2023).
-Given these limitations, recent research has begun exploring multi-agent systems to enhance vulnerability detection (Wen et al., 2024; Hu et al., 2023; Wei et al., 2024).
-Multi-agent approaches distribute specific tasks among role-specific agents while allowing discussion and more information exchange between agents.
-GPTLens (Hu et al., 2023) is one such approach, dividing vulnerability detection into two stages, generation by an auditor agent and discrimination by a critic agent.
-In GPTLens, the final decision is made by the critic agent, which is tasked to be a harsh vulnerability critic and doubt the vulnerabilities and reasoning given by the auditor agent.
+Single-agent systems are inherently constrained by their reliance on a single perspective when analyzing vulnerabilities, which can lead to issues such as hallucination and overconfidence in incorrect reasoning (zhang2023siren; shi2025mitigating; huang2025survey; liang2023encouraging; du2023improving).
+These single-agent systems may fail to critically verify their own output, which is crucial when identifying subtle vulnerabilities (liang2023encouraging).
+Given these limitations, recent research has begun exploring multi-agent systems to automate software engineering tasks (he2025llm; hu2023large; wei2024llm).
+Such approaches distribute tasks across role-specific agents while allowing discussion and richer information exchange between agents.
 In this work, we propose VulTrial, a novel multi-agent LLM framework that employs a courtroom-inspired scenario.
 VulTrial features four role-specific agents: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury).
-Unlike GPTLens, where the agent that does the critic delivers the final result, VulTrial delegates the final prediction to a dedicated review board agent, which is designed to be unbiased toward the security researcher and code author agents.
 The review board bases its decision on arguments from the security researcher (highlighting potential vulnerabilities), the code author (defending the code), and the moderator (summarizing debate points).
-Review board agent works similarly to the jury in the courtroom scenario, where they need to hear all the evidence presented in the case and make a final verdict based on it.
-Additionally, VulTrial allows for customizable rounds of interaction among the first three agents, which can provide opportunities for rebuttal and refinements.
-We evaluate VulTrial results using two base LLMs, GPT-3.5 and GPT-4o (see Section 3.3).
-We benchmarked VulTrial against single-agent baselines from PrimeVul (Ding et al., 2025b) and a multi-agent baseline, GPTLens (Hu et al., 2023), all of which also utilize the same base LLMs for a fair comparison.
-PrimeVul (Ding et al., 2025b) (chain of thought and fine-tuning) and a multi-agent baseline, GPTLens (Hu et al., 2023).
-In the experiment using GPT-3.5, VulTrial outperforms the single-agent and multi-agent baselines by 151.69% and 239.78% respectively.
-Similarly, using GPT-4o, VulTrial also improves the performance of single-agent and multi-agent baselines by 102.39% and 84.17%.
-Furthermore, to enhance VulTrial’s performance and analyze the impact of instruction tuning on a role-specific agent, we conducted additional experiments by applying instruction tuning to each agent in VulTrial.
-For the instruction tuning dataset, we selected 50 sample pairs from the validation data.
-The results indicate that tuning the moderator agent in VulTrial produced the most significant improvements, achieving an 18.53% increase compared to its non-tuned counterpart and outperforming the single-agent and multi-agent baselines by 139.89% and 118.30%, respectively.
-We also examined the effect of increasing the number of discussion rounds in VulTrial and observed that, when using GPT-4o, adding more rounds does not result in improved performance.
-Finally, an ablation study revealed that each agent within VulTrial is crucial for achieving the best overall performance.
+The review board agent works similarly to a jury in the courtroom scenario, where they need to hear all the evidence presented in the case and make a final verdict based on it.
+In VulTrial, we use the same underlying LLM for all agents, building on prior works (wei2022chain; chatterjee2024posix; hu2023large) that show prompt-induced roles can lead to different reasoning paths using the same model.
+We benchmarked VulTrial against single-agent and multi-agent baselines.
+For the single-agent, we include deep learning (DL)-based models (CodeBERT (feng2020codebert), CodeT5 (wang2021codet5), UniXCoder (guo2022unixcoder), and LineVul (fu2022linevul)), as well as an LLM-based method (Ding et al.’s CoT prompts (ding2024vulnerability)).
+For the multi-agent setting, we utilized GPTLens (gptlens).
+All LLM-based methods are instantiated using GPT-3.5 and GPT-4o as the base model (see Section 3.3).
+The results show that VulTrial improves over the baselines.
+With GPT-4o, VulTrial nearly doubles the performance of both the best performing single-agent and multi-agent baselines (P-C score=81 vs.
+40 and 44, respectively).
+To assess the effect of instruction tuning on role-specific agents, we tuned each agent in VulTrial using 50 validation sample pairs.
+Instruction tuning the moderator agent in VulTrial improves the P-C score by 15 (18.52%) over the non-tuned version and surpasses the single-agent and multi-agent baselines by 56 (140%) and 52 (118%).
+To assess generalizability, we additionally confirmed the efficacy of VulTrial using LLaMA-3.1-8B, a publicly available model that can be deployed in-house.
+Separately, we perform an ablation study to verify that each agent contributes positively to achieving the best performance.
+Beyond quantitative results, we also qualitatively evaluate the generated explanations, examining completeness, clarity, actionability, and informativeness, and observe strong performance across all four dimensions.
+Finally, we deploy VulTrial in the wild, where it successfully uncovers multiple confirmed zero-day vulnerabilities, demonstrating its practical effectiveness.
 To summarize, the contributions of our study are as follows:
   • We propose VulTrial, a novel multi-agent, courtroom-inspired approach leveraging LLMs for automated vulnerability detection.
-VulTrial consists of role-specific agents (i.e., security researcher, code author, moderator, and review board).
-The ablation study highlights that each agent in VulTrial is necessary to get the best performance.
-  • We demonstrate that targeted instruction tuning of role-specific agents in VulTrial improves vulnerability detection performance compared to the baselines.
-Even with limited training examples (as few as 50 pairs), we get an improvement of 18.53% compared to the non-tuned version.
-  • We conduct extensive experiments using GPT-3.5 and GPT-4o in PrimeVul dataset (Ding et al., 2025b) and compare the VulTrial with single-agent and multi-agent baselines.
-Overall, the results using VulTrial outperform the baselines.
+Also, different from many prior approaches that only output a binary label, VulTrial provides detailed explanations for each classification.
+  • We demonstrate that targeted instruction tuning of role-specific agents in VulTrial improves the performance.
+  • We conduct extensive experiments demonstrating the efficacy of VulTrial on benchmark data and in the wild.
 2. VulTrial
-Our proposed method, namely VulTrial, works by utilizing multiple agents, all of which are powered by LLMs.
-VulTrial consists of four role-specific agents, each inspired by distinct roles in a courtroom setting: a security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury).
-This structured debate encourages thorough, adversarial discussions, allowing for deeper analyses and refined conclusions regarding code vulnerabilities.
-Drawing inspiration from courtroom practices, where structured adversarial processes have been shown to reduce wrongful convictions and improve the fairness of judgments (Findley and Scott, 2006; Garrett, 2011), VulTrial aims to improve vulnerability detection in a similar manner.
-The overall architecture of these agents is illustrated in Fig.
+VulTrial leverages multiple LLM-based agents modeled after courtroom roles: a security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury).
+This structured debate enables thorough, adversarial discussions for deeper analyses and refined conclusions regarding code vulnerabilities.
+Drawing inspiration from courtroom practices, where structured adversarial processes have been shown to reduce wrongful convictions and improve the fairness of judgments (findley2006multiple; garrett2011convicting), VulTrial aims to improve vulnerability detection in a similar manner.
+VulTrial architecture is shown in Fig.
 1, with an example output in Fig.
@@
 The third agent, which is the moderator agent, is tasked with summarizing the arguments presented by both the security researcher and the code author agents to ensure a structured exchange of viewpoints between them.
-VulTrial supports multiple rounds of interaction among the first three agents before passing their viewpoints into the last agent.
+VulTrial supports multiple rounds of interaction among the first three agents before passing their viewpoints to the last agent.
 The last agent, namely the review board agent, is tasked with analyzing all the information given by the moderator, the security researcher, and the code author agents.
 Based on this collective information, the review board delivers a final verdict regarding the security status of the code.
-This verdict highlights whether the potential vulnerability brought up by the security researcher agent is indeed a vulnerability.
-We chose to have the agents simulate a jury sentencing scenario over a judge sentencing scenario, as previous studies have shown that jury participation can help prevent excessively harsh punishments (Hans et al., 2015; McLeod, 2024).
-The details of each agent and their interaction will be discussed in the following subsections.
+We chose to have the agents simulate a jury over a judge sentencing scenario, as previous studies have shown that jury participation can help prevent excessively harsh punishments (hans2015death; mcleod2024democratic).
+The details of the agent’s interaction will be discussed in the following subsections.
+The prompts can be found in our online appendix (ourreplicationpackageall).
 Figure 1.
-Overview of the VulTrial architecture.
-It employs a multi-agent framework inspired by a courtroom setting to assess code vulnerabilities.
-The security researcher (prosecutor) identifies potential vulnerabilities, the code author (defense attorney) provides counterarguments, the moderator (judge) summarizes key points, and the review board (jury) delivers the final verdict.
+VulTrial architecture which uses a courtroom-inspired multi-agent framework where the security researcher (prosecutor) identifies potential vulnerabilities, the code author (defense attorney) provides counterarguments, the moderator (judge) summarizes key points, and the review board (jury) delivers the final verdict on code vulnerabilities.
 2.1. Security Researcher Agent (Prosecutor Role)
@@
 In the case of vulnerability detection, the security researcher agent inspects the provided code, identifies potential vulnerabilities, and presents a rationale for why these vulnerabilities may pose a threat.
-An example of the security researcher agent’s output is shown in Fig.
+An example is shown in Fig.
 2, where it highlights two potential vulnerabilities in the code, along with their reasoning and impact.
-The prompt used for this agent is highlighted in Fig.
-3.
-In the first round, the agent only receives a code function and determines whether the code is vulnerable.
-In subsequent rounds, it also receives the output generated by two other agents (i.e., code author and moderator) which is highlighted as ¡chat history¿ in Fig.
-3.
-Beyond simply reporting any vulnerability, the agent is explicitly asked to include (1) a description of each vulnerability, (2) the reasoning behind its identification, and (3) the potential impact if that vulnerability were to be exploited.
-We asked for this agent to include the reasoning as it has been highlighted in previous studies (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024) that prompting LLMs to articulate their reasoning processes can significantly enhance their performance.
-Furthermore, including the potential impact of each vulnerability gives the other agents (i.e., the code author, moderator, and review board) a clearer perspective when they form their answer.
+In the first round, the agent receives a code function and determines whether the code is vulnerable.
+In subsequent rounds, it also receives the output generated by two other agents (i.e., code author and moderator).
+Beyond simply reporting any vulnerability, the agent is explicitly asked to include (1) a description of each vulnerability, (2) the reasoning behind it, and (3) the potential impact if exploited.
+We require the agent to include its reasoning, as previous studies (wei2022chain; kojima2022large; cheng2024chainlm) show that prompting LLMs to articulate their reasoning can significantly enhance their performance.
+Furthermore, including the potential impact of each vulnerability gives the other agents a clearer perspective when they form their answer.
+2.2. Code Author Agent (Defense Attorney Role)
+The code author agent in VulTrial acts similarly to a defense attorney in a courtroom scenario.
+While the prosecutor aims to prove that the defendant is guilty, the defense attorney counters these allegations by advocating for the defendant’s innocence.
+Applying this scenario to vulnerability detection, the code author agent seeks to defend the integrity and security of their code, providing arguments and evidence to counter the security researcher agent’s claims.
+If the vulnerability is valid, the agent can acknowledge the issue and propose effective mitigation strategies.
+The prompt provides the agent with the code snippet under investigation, along with information generated by the security researcher agent, which outlines a list of potential vulnerabilities (ourreplicationpackageall).
+The agent is then tasked with clearly articulating the following outputs: (1) identification of the vulnerability being discussed, (2) explicit acknowledgment or refutation of the vulnerability, and (3) detailed reasoning justifying the refutation or mitigation methods in case the vulnerability is acknowledged.
+An example of the code author agent’s output is also shown in Fig.
+2, where it refutes one vulnerability and provides reasoning for it.
+Since the security researcher agent may highlight multiple potential vulnerabilities, it is crucial for the code author agent to clearly specify which vulnerability it addresses in each response.
+Explanations from the code author agent enable other agents (the moderator and review board) to better understand the justification or rationale behind the agent’s position.
+For the security researcher agent, this reasoning would be used for building their rebuttal.
+This structured approach helps ensure that subsequent agents are equipped to offer well-informed counterarguments, summaries, or final evaluations.
+2.3. Moderator Agent (Judge Role)
+Just as judges provide guidance to jurors and uphold fairness in a courtroom scenario, within the context of vulnerability detection, this agent is tasked with objectively summarizing the arguments presented by both the security researcher agent (prosecutor) and the code author agent (defense attorney).
+It should focus only on the facts, avoiding any bias toward either side.
+This impartial summary helps the review board agent in reaching a decision, while also clarifying critical discussion points for subsequent rounds of argumentation by the security researcher and code author agents.
+In addition to the code snippet, the moderator receives the security researcher’s identified vulnerabilities and the code author’s counterarguments.
+The moderator’s output consists of concise summaries of these two viewpoints, highlighting the key points from each.
+Fig.
+2 shows an example of the moderator agent summarizing the arguments from the security researcher (e.g., the number of potential vulnerabilities identified, their main reasons, and impacts) as well as from the code author, such as which vulnerabilities were acknowledged or refuted with the main reasoning.
 Figure 2.
@@
 The code is vulnerable as it does not check for negative indices, allowing an attacker to read arbitrary data by crafting a model with negative values in indices.
-You are the Security Researcher.
-Identify all potential security vulnerabilities in the given code snippet.
-Provide your output as a JSON array. Each element in the array represents one identified vulnerability and should include:
-vulnerability”: A short name or description of the vulnerability.
-reason”: A detailed explanation of why this is a vulnerability and how it could be exploited.
-impact”: The potential consequences if this vulnerability were exploited.
-¡code¿
-¡chat history¿
-Figure 3. Security researcher agent prompt.
-2.2. Code Author Agent (Defense Attorney Role)
-The code author agent in VulTrial acts similarly to a defense attorney in a courtroom scenario.
-While the prosecutor aims to prove that the defendant is guilty, the defense attorney counters these allegations by advocating for the defendant’s innocence.
-Applying this scenario to the vulnerability detection, the code author agent seeks to defend the integrity and security of their code, providing arguments and evidence to counter the security researcher agent’s claims.
-If the vulnerability is valid, the agent can acknowledge the issue and propose effective mitigation strategies.
-The specific prompt for the code author agent is shown in Fig.
-4.
-This prompt provides the agent with the code snippet under investigation along with information generated by the security researcher agent (included in ¡chat history¿), which outlines a list of potential vulnerabilities.
-The agent is then tasked with clearly articulating the following outputs: (1) identification of the vulnerability being discussed, (2) explicit acknowledgment or refutation of the vulnerability, and (3) detailed reasoning justifying the refutation or, in cases where the agent acknowledges the vulnerability, proposed methods for mitigation.
-An example of the code author agent’s output is also shown in Fig.
-2, where it acknowledges one of the vulnerabilities identified by the security researcher and proposes a mitigation strategy.
-Additionally, it refutes the other vulnerability and provides reasoning for it.
-Since the security researcher agent may highlight multiple potential vulnerabilities, it is crucial for the code author agent to clearly specify which vulnerability it addresses in each response.
-As emphasized by prior studies (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024), explicitly prompting LLMs to articulate their reasoning has shown significant improvements in the accuracy and reliability of their outputs.
-Furthermore, clear explanations from the code author agent enable other agents (the moderator and review board) to better understand the justification or rationale behind the agent’s position.
-For the security researcher agent, this reasoning would be used for building their rebuttal response.
-This structured approach helps ensure that subsequent agents are equipped to offer well-informed counterarguments, summaries, or final evaluations.
-You are the Code Author of the attached code.
-The Security Researcher has presented a JSON array of alleged vulnerabilities.
-You must respond as if you are presenting your case to a group of decision-makers who will evaluate each claim.
-Your tone should be respectful, authoritative, and confident, as if you are defending the integrity of your work to a panel of experts.
-For each identified vulnerability, produce a corresponding JSON object with the following fields:
-vulnerability”: The same name/description from the Security Researcher’s entry.
-response-type”: ‘refutation’ if you believe this concern is unfounded, or ‘mitigation’ if you acknowledge it and propose a workable solution.
-reason”: A concise explanation of why the vulnerability is refuted or how you propose to mitigate it.
-¡code¿
-¡chat history¿
-Figure 4. Code author agent prompt.
-2.3. Moderator Agent (Judge Role)
-In VulTrial, the moderator agent attempts to mirror the role of a judge in a courtroom scenario.
-Just as judges provide guidance to jurors and uphold fairness, within the context of vulnerability detection, this agent is tasked with objectively summarizing the arguments presented by both the security researcher agent (prosecutor) and the code author agent (defense attorney).
-It should focus only on the facts, avoiding any bias toward either side’s party.
-The primary purpose of this impartial summary is to assist the review board agent in reaching a decision, while also clarifying critical discussion points for subsequent rounds of argumentation by the security researcher and code author agents.
-The prompt for this agent is highlighted in Fig.
-5.
-In addition to accessing the code snippet in question, the moderator receives the security researcher’s identified vulnerabilities and the code author’s counterarguments (included in ¡code history¿).
-The moderator’s output consists of concise summaries of these two viewpoints, highlighting the key points from each.
-Fig.
-2 presents an example of the moderator agent’s output.
-It summarizes the arguments from the security researcher (e.g., the number of potential vulnerabilities identified, their main reasons, and impacts) as well as from the code author, such as which vulnerabilities were acknowledged or refuted with the main reasoning.
-You are the Moderator, and your role is to provide a neutral summary.
-After reviewing both the Security Researcher’s identified vulnerabilities and the Code Author’s responses, provide a JSON object with two fields:
-security researcher summary”: A concise summary of the vulnerabilities and reasoning presented by the Security Researcher.
-author summary”: A concise summary of the Code Author’s counterarguments or mitigation strategies.
-¡code¿
-¡chat history¿
-Figure 5. Moderator agent prompt.
 2.4. Review Board (Jury Role)
-The review board agent in this VulTrial is similar to the role of jury in a courtroom scenario.
 In a traditional courtroom, the jury weighs all evidence and testimony presented during a trial, alongside any guidance from the judge, to determine whether a defendant is guilty or not.
@@
 Its purpose is to deliver a final verdict on whether the potential vulnerabilities identified by the security researcher are indeed valid.
-The prompt guiding the review board agent is shown in Figure 6.
 This agent receives (1) the summary from the moderator agent; (2) the list of potential vulnerabilities from the security researcher agent; (3) the code author agent’s counterarguments; (4) the original code snippet.
@@
 First, it must extract the names of potential vulnerabilities from the list provided by the security researcher and proceed with a decision for each one.
-The review board agent needs to highlight the specific vulnerability being discussed as there may be multiple vulnerabilities under review.
 Second, the agent must determine whether each identified vulnerability is “valid,” “invalid,” or “partially valid.” This process is similar to a jury deciding whether a defendant is guilty or not guilty.
 Third, the agent must assess the severity of the vulnerability, similar to how a jury determines the extent of damages in a trial.
-However, instead of assigning an open-ended value, the review board needs to categorize severity as “low,” “medium,” “high.” If the vulnerability is deemed invalid, the severity is marked as “none.” Fourth, the agent must provide a recommended action based on the decision and severity level.
+However, instead of assigning an open-ended value, the review board needs to categorize severity as “low,” “medium,” or “high.” If the vulnerability is deemed invalid, the severity is marked as “none.” Fourth, the agent must provide a recommended action based on the decision and severity level.
 For example, if a vulnerability is deemed valid and has a high severity, the recommended action should be “fix immediately.” Finally, the agent must generate an explanation for its decisions.
-Unlike a jury, which is not required to provide reasoning, we ask the review board agent to do so, as asking LLMs for reasoning can enhance performance (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024) and helps users to understand the decision.
-We employ a threshold-based approach to classify a function as vulnerable, rather than a simple binary “vulnerable or not,” because it offers flexibility for users to tailor vulnerability detection to their tolerance for risk and false positives.
-By adjusting the threshold, users can decide how strict they want the classification to be.
-For instance, if a user prefers to catch even the smallest chance of vulnerability (and is comfortable with a higher rate of false positives), they can opt to flag any partially valid vulnerabilities regardless of severity.
-Conversely, if a user wants fewer false positives, they can raise the threshold to consider only higher-severity issues as triggers.
-In our study, we flag a function as vulnerable if at least one vulnerability raised by the security researcher agent is deemed valid at high severity with recommended action fix immediately by the review board agent.
-This stricter threshold ensures that only the critical and well-substantiated vulnerabilities contribute to the classification, reducing false positives from lower-risk or ambiguous issues.
+Unlike a real jury, we require the review board agent to provide reasoning, as it improves LLMs’ performance (wei2022chain; kojima2022large; cheng2024chainlm) and helps users understand the decision.
+We use a threshold rather than a simple binary “vulnerable or not,” because it offers flexibility for users to tailor vulnerability detection to their tolerance for risk and false positives.
+For instance, if a user prefers to catch even the smallest chance of vulnerability, they can opt to flag any partially valid vulnerabilities.
+Conversely, if a user wants fewer false positives, they can raise the threshold to consider only higher-severity issues.
+We flagged a function as vulnerable if at least one vulnerability raised by the security researcher agent is deemed valid at high severity with recommended action fix immediately by the review board agent.
+This threshold ensures that only the critical and well-substantiated vulnerabilities contribute to the classification, reducing false positives from lower-risk or ambiguous issues.
 We aim to reduce noise by excluding vulnerabilities marked as partially valid or those with lower severity levels.
@@
 Meanwhile, the second vulnerability (Type Handling) is marked as “partially valid” with “medium” severity.
-Since at least one confirmed vulnerability is present, the final verdict for this code is vulnerable.
-You are the Review Board.
-After reviewing the Moderator’s summary, Code Author’s and Security Researcher’s argument, and ¡code¿, produce a JSON array of verdicts for each vulnerability identified by the Security Researcher.
-Each object in the array should include: vulnerability”: The same name as given by the Security Researcher.
-decision”: One of ‘valid’, ‘invalid’, or ‘partially valid’.
-severity”: If valid or partially valid, assign a severity (‘low’, ‘medium’, ‘high’); if invalid, use ‘none’.
-recommended action”: Suggest what should be done next (e.g., ‘fix immediately’, ‘monitor’, ‘no action needed’).
-reason”: A brief explanation of why you reached this conclusion, considering both the Security Researcher’s and Code Author’s perspectives.
-¡code¿
-¡chat history¿
-Figure 6. Review board agent prompt.
+As at least one confirmed vulnerability is present, the final verdict for this code is vulnerable.
 2.5. Agent Interaction
 In VulTrial, four agents interact sequentially: the security researcher agent, the code author agent, the moderator agent, and finally the review board agent.
-During the first round, the security researcher agent (1) only has access to the original code snippet.
+During the first round, the security researcher agent (1) has access only to the original code snippet.
 In subsequent rounds, it also receives information from the code author and moderator.
-Similarly, the code author agent (2) initially sees only the security researcher agent’s findings, but in the next rounds it can also incorporate the moderator’s feedback.
+Similarly, the code author agent (2) initially sees only the security researcher agent’s findings, but in the next rounds, it can also incorporate the moderator’s (3) output.
 This setup is similar to a courtroom process, which can allow for multiple rounds of arguments and counterarguments between the prosecutor, defense attorney, and judge.
@@
 Allowing multiple discussion rounds may enhance the iterative refinement of analyses and arguments.
-It would also provide more chances for the security researcher agent and code author agent for rebuttal.
-Nevertheless, increased rounds entail higher operational costs.
-Therefore, it is essential to balance the trade-off between performance improvement and the cost.
-In this study, we initially limit the number of discussion rounds to one (see Section 4.1), and subsequently, in Section 4.3, we explore how varying the number of discussion rounds impacts the overall performance of VulTrial.
+It also provides both the security researcher and code author agents with more opportunities for rebuttal.
+In this study, we initially limit the number of discussion rounds to one (Section 4.1) and then assess how changing the number of rounds affects VulTrial’s performance (Section 4.3).
 3. Experimental Setup
 3.1. Dataset
-In our study, we utilized the PrimeVul dataset, recently proposed by Ding et al.
-(Ding et al., 2025b), to evaluate LLMs using single-agent and multi-agents on vulnerability detection.
-Specifically, we used the PrimeVul pair data, which presents a challenging benchmark, as the vulnerable and benign code functions share at least 80% of their string content.
-This dataset ensures no data leakage or exact duplication between the training, testing, and validation sets.
-The latest version of PrimeVul Pair contains 435 vulnerable and 435 benign functions (Ding et al., 2025a).
+We utilized the PrimeVul dataset, recently proposed by Ding et al.
+(ding2024vulnerability), to evaluate vulnerability detection approaches.
+Specifically, we used the PrimeVul pair dataset, which presents a challenging benchmark, as the vulnerable and benign code functions share at least 80% of their string content.
+The pair dataset includes 3,789 training pairs, 480 validation pairs, and 435 test pairs, with no data leakage across splits.
+In contrast, the full PrimeVul dataset contains 175,797 training functions (4,862 vulnerable and 170,935 benign) and 23,948 validation functions (593 vulnerable and 23,355 benign).
 3.2. Research Questions
@@
 RQ1: How does the performance of VulTrial compare to that of single-agent and multi-agent baselines on PrimeVul?
-Prior work (Ding et al., 2025b) demonstrated the poor performance of GPT-3.5 and GPT-4 for vulnerability detection on pairwise vulnerable and benign functions in PrimeVul.
-These findings highlight the limitations of single-agent prompting strategies in challenging detection scenarios.
-In this RQ, we investigate whether VulTrial, a multi-agent LLM-based framework (Section 2), can outperform both single-agent methods (Ding et al., 2025b) and existing multi-agent approaches such as GPTLens (Hu et al., 2023) on PrimeVul.
+Prior work (ding2024vulnerability) demonstrated the poor performance of GPT-3.5 and GPT-4 for vulnerability detection on pairwise vulnerable and benign functions in PrimeVul.
+These findings highlight the limitations of single-agent prompting strategies for this challenging setting.
+In this RQ, we investigate whether VulTrial, a multi-agent LLM-based framework (Section 2), can outperform both single-agent methods (ding2024vulnerability) and existing multi-agent approaches such as GPTLens (hu2023large) on the PrimeVul pair.
 RQ2: How does instruction tuning on role-specific agents influence the performance of automated vulnerability detection?
@@
 RQ3: Does allowing multiple rounds of debate among role-specific agents in VulTrial lead to improvement compared to a single-pass analysis?
-In previous RQs, we focus on conducting a single round of discussions among agents to reach a final verdict.
+In previous RQs, we focused on conducting a single round of discussions among agents to reach a final verdict.
 However, allowing multiple rounds of discussion could give agents the opportunity to refine their arguments to reach a more accurate decision.
@@
 While the results from previous RQs highlight VulTrial’s effectiveness, it is not known whether each agent is essential to the final prediction and how it impacts the performance.
-In this RQ, we do the ablation study that systematically omits one agent at a time to help us verify each agent’s contribution and evaluate whether the loss of a role reduces the performance.
-Important note: The *ground-truth label* for this code is: “¡vulnerable/benign¿ vulnerable” means this code contains at least one vulnerability.
-benign” means there are no vulnerabilities in this code.
-Use the label to ensure your arguments match the fact: Even if you are “defending” the code, if the label is “vulnerable,” you should not fully deny actual issues.
-If the label is “benign,” you can confidently refute all vulnerability claims.
-Here is the CVE ID and description that is related with the code:
-¡CVE ID¿”: ¡CVE Description¿
-Figure 7. Example of additional information provided to the code author agent for constructing the instruction tuning dataset.
+In this RQ, we do the ablation study that omits one agent at a time to verify each agent’s contribution.
 3.3. Baseline and Models
 Baselines.
-For the single-agent baseline, we utilized the chain of thought (CoT) (Wei et al., 2022) prompting technique, which achieved the highest results in the PrimeVul (Ding et al., 2025b).
+We employed both DL-based and LLM-based vulnerability detection techniques as single-agent baselines.
+For the DL-based methods, we used several state-of-the-art models widely adopted in prior vulnerability detection studies (ding2024vulnerability; xiong2024vuld; zhou2024large; chen2023diversevul):
+  • CodeBERT (feng2020codebert): A bimodal pre-trained transformer model trained on source code and natural language.
+  • CodeT5 (wang2021codet5): A transformer model that learns to understand and generate code across multiple programming languages.
+  • UniXCoder (guo2022unixcoder): A cross-modal pre-trained model that jointly learns representations of code, ASTs, and execution traces.
+  • LineVul (fu2022linevul): A line-level transformer-based vulnerability prediction model detecting vulnerable lines through fine-grained context representations.
+For CodeBERT, CodeT5, and UniXCoder, we performed two separate fine-tuning: (1) using the full PrimeVul training and validation sets, and (2) using only the pair-wise subsets from PrimeVul.
+For LineVul, we first evaluated the model provided in the previous study (fu2022linevul).
+In addition, we further fine-tuned LineVul using the pair-wise subset to ensure a consistent comparison with the other DL models.
+For the LLM-based approach, we utilized the chain of thought (CoT) (wei2022chain) prompting technique, which achieved the highest results in the PrimeVul (ding2024vulnerability).
 CoT prompting encourages LLMs to reason through problems step by step to arrive at an answer.
-For the multi-agent baseline, we utilized an approach from a recent previous study, GPTLens (Hu et al., 2023).
+For the multi-agent baseline, we utilized an approach from a recent study, GPTLens (hu2023large).
 GPTLens operates with two types of agents: auditor agent, which identifies potential vulnerabilities, and a critic agent, which evaluates and scores the identified vulnerabilities.
@@
 Unlike GPTLens, where the critic agent determines the final decision, VulTrial used a different agent (i.e., review board agent) for this task, designed to be neutral toward the security researcher and code author agents.
-Recent studies (Gao et al., 2024; Hu et al., 2024) suggest that employing domain-specific multi-agent systems that reflect a more realistic scenario leads to better decision making, further supporting our approach in VulTrial.
-Models.
-We employed multiple models in this study.
-First, we used several well-established OpenAI models known for their strong performance, as confirmed by prior research (Widyasari et al., 2024; Hu et al., 2023; Zhou et al., 2024).
-Specifically, we utilized GPT-3.5 and GPT-4o.
-GPT-3.5 (OpenAI, 2023a): GPT-3.5 is one of the first widely popular OpenAI LLM models.
-Following PrimeVul (Ding et al., 2025b), we also evaluate the performance of this model in our study.
+Recent studies (gao2024large; hu2024unlocking) suggest that employing domain-specific multi-agent systems that reflect a more realistic scenario leads to better decision making, further supporting our approach in VulTrial.
+Models. We used well-established OpenAI models known for their strong performance, as confirmed by prior research (widyasari2024beyond; hu2023large; zhou2024large_emerging).
+GPT-3.5 (openai2023gpt35): GPT-3.5 is one of the first widely popular OpenAI LLM models.
+Following PrimeVul (ding2024vulnerability), we also evaluate the performance of this model in our study.
 In our evaluation, we utilized gpt-3.5-turbo-0125, the latest iteration of the GPT 3.5 model with a context window of 16k tokens.
-GPT-4o (OpenAI, 2024b): We employed GPT-4o, one of OpenAI’s flagship large language models recommended as a replacement for GPT-4 (OpenAI, 2023b).
-We chose GPT-4o over GPT-4 (used in PrimeVul’s evaluation) because OpenAI indicates that GPT-4o offers performance comparable to GPT-4 while significantly reducing computational costs (OpenAI, 2024b).
+GPT-4o (openai2024gpt4o): We chose GPT-4o, one of OpenAI’s flagship models, over GPT-4 (openai2023gpt4) (used in PrimeVul’s evaluation) because OpenAI indicates that GPT-4o offers performance comparable to GPT-4 while significantly reducing computational costs (openai2024gpt4o).
 Specifically, we used gpt-4o-2024-08-06 allows up to 128k tokens in its context window and 16k output tokens.
 3.4. Evaluation Metrics
-Following the approach in PrimeVul (Ding et al., 2025b), we adopt a pair-wise evaluation strategy.
+Following the approach in PrimeVul (ding2024vulnerability), we adopt a pair-wise evaluation strategy.
 It is designed to assess the model’s prediction on pairs of vulnerable and benign code, rather than the single codes.
 This method highlights the model’s capacity to differentiate subtle variations that make one function vulnerable and the other benign.
-The four key metrics are defined as follows:
-  • Pair-wise Correct Prediction (P-C): The model correctly predicted both codes in the pair.
-  • Pair-wise Vulnerable Prediction (P-V): The model incorrectly classifies both code as vulnerable.
-  • Pair-wise Benign Prediction (P-B): The model incorrectly classifies both code as benign.
-  • Pair-wise Reversed Prediction (P-R): The model assigns the opposite label to each code in the pair, predicting “benign” for a vulnerable function and “vulnerable” for a benign one.
-We use P-C as our primary evaluation metric. A high P-C indicates accurate vulnerability detection that avoids false positives and negatives, making it a more robust indicator of overall performance.
+We report five metrics.
+Pair-wise Correct Prediction (P-C) records pairs where both functions are classified correctly.
+Pair-wise Vulnerable Prediction (P-V) records pairs where both functions are incorrectly labeled as vulnerable, while Pair-wise Benign Prediction (P-B) records pairs labeled incorrectly as benign.
+Pair-wise Reversed Prediction (P-R) captures cases where the model assigns opposite labels to the pair, predicting “benign” for the “vulnerable” code and “vulnerable” for the “benign” one.
+Pair-wise Error (E) reflects pairs in which one function is misclassified.
+Among these metrics, P-C is our primary evaluation metric.
+A high P-C indicates accurate vulnerability detection that avoids false positives and negatives, making it a more robust indicator of overall performance.
+To provide more comprehensive discussion in VulTrial and baselines comparison, we also report standard classification metrics: Precision (P), Recall (R), and False Positive Rate (FPR), following the formula used in the previous study (ding2024vulnerability).
+These traditional metrics provide additional insights into the model’s general classification behavior, particularly in individual predictions.
 3.5. Methods
 3.5.1. Answering RQ1:
-We compare VulTrial with both single-agent (Ding et al.’s CoT(Ding et al., 2025b)) and multi-agent (GPTLens) baselines, using two different LLMs, i.e., GPT-3.5 and GPT-4o.
-Section 3.3 provides further details on our baseline methods and model choices.
-We evaluate each approach using several metrics, with the Pair-wise Correct (PC) rate as our primary evaluation metric.
-A higher PC rate indicates better performance.
+We compare VulTrial with both single-agent (CodeBERT (feng2020codebert), CodeT5 (wang2021codet5), UniXCoder (guo2022unixcoder), LineVul (fu2022linevul), and Ding et al.’s CoT(ding2024vulnerability)) and multi-agent (GPTLens) baselines.
+For the LLM-based techniques, we utilize two different LLMs, i.e., GPT-3.5 and GPT-4o (see Section 3.3).
+We evaluate each approach using several metrics (P-C, P-V, P-B, P-R, Error, P, R, and FPR).
 3.5.2. Answering RQ2:
 For this RQ, we conducted instruction tuning using samples drawn from GPT-4o’s validation set.
-After running GPT-4o on the validation set, we took its correct outputs to create an instruction tuning dataset for a GPT-4o model.
-We selected GPT-4o because, as shown in RQ1, VulTrial using GPT-4o outperforms both single-agent and multi-agent baselines and achieves the best overall results, making it an ideal candidate for evaluating the impact of instruction tuning.
+After running GPT-4o on the set, we collected its correct outputs to build an instruction-tuning dataset for a GPT-4o model.
+We focus on GPT-4o because, as shown in RQ1, VulTrial with GPT-4o outperforms both single-agent and multi-agent baselines and achieves the best overall results.
 Our goal is to determine whether instruction tuning can further enhance its performance.
@@
 For instruction tuning data, we use 50 code pairs (vulnerable and benign) with the fewest tokens in the validation set to help manage the tuning cost.
-In addition, OpenAI has highlighted that GPT-4o can achieve strong results with “as few as a few dozen examples” (OpenAI, 2024a).
-When creating our instruction tuning dataset, we initially included the true label in the prompt (e.g., adding vulnerability-related information for the code author agent, see Fig.
-7), but then removed direct label disclosures in the final dataset.
+We chose 50 samples based on preliminary experiments comparing different dataset sizes (25, 50, and 75 pairs), as 50 resulted in significant performance gains over 25 samples, while achieving comparable results to using 75 samples (ourreplicationpackageall).
+In addition, OpenAI has highlighted that GPT-4o can achieve strong results with “as few as a few dozen examples” (openai2024gpt4o_finetuning).
+When creating our instruction tuning dataset, we initially included the true label in the prompt, but then removed direct label disclosures in the final dataset.
+For example, for the code author agent (ourreplicationpackageall), we included vulnerability-related information such as CVE ID and label.
+Similarly, for the security researcher agent, we added the CVE description and label for both vulnerable and benign code, prompting the agent to connect code to relevant vulnerabilities.
+For the review board agent, the prompt included ground-truth labels but also emphasized that the agent should not reference them explicitly in its reasoning.
+In contrast, the moderator agent’s prompt excluded any ground-truth vulnerability information and was cleaned to avoid indirect label leakage from chat history.
 As a result, each prompt retained its original role-specific structure and was paired with the correct assistant output.
-We utilized following scenarios for the evaluation: (1) no instruction tuning, (2) instruction tuning only the security researcher agent, (3) instruction tuning only the code author agent, (4) instruction tuning only the moderator agent, (5) instruction tuning only the review board agent, (6) instruction tuning both security researcher and code author agents, (7) instruction tuning the security researcher, code author, and moderator agents, and (8) instruction tuning all agents.
+For this instruction tuning data, we ensure that the agent’s predicted label (“vulnerable” or “benign”) matches the ground-truth label.
+Additionally, we manually verify the quality of the agent’s explanation, confirming that it correctly identifies the presence or absence of a vulnerability.
+An example of the vulnerable instance is shown in the online appendix (ourreplicationpackageall); this vulnerability is classified as CWE-416: Use After Free.
+The security researcher agent correctly identifies the issue within the CleanWriters function, where the sequence of operations may lead to undefined behavior if gf_isom_box_del() uses the writer object after it has been freed.
+The explanation aligns well with the root cause described in CVE-2020-35980.
+For the corresponding benign variant (ourreplicationpackageall), the code author agent correctly refutes the presence of a vulnerability by noting that gf_list_del_item() only removes the items from the list without deallocating their memory.
+The justification provided by the review board agent also matches the code author’s agent clarification and supports the benign ground-truth label.
+We evaluated eight scenarios: (1) no instruction tuning; (2-5) tuning only the security researcher, code author, moderator, or review board agent; (6) tuning security researcher and code author; (7) tuning security researcher, code author, and moderator; and (8) tuning all agents.
 To systematically assess the impact of instruction tuning, we first applied it to individual agents to understand their specific contributions.
 Then, we incrementally introduced instruction tuning in the order of agent interaction, allowing us to analyze how tuning at different stages influences the overall performance of VulTrial.
-This stepwise approach provides insights into the cumulative effects of instruction tuning across multi-agents.
+This stepwise approach provides insights into the cumulative effects of instruction tuning across multiple agents.
 3.5.3. Answering RQ3:
@@
 We selected these values to systematically investigate the impact of multiple interaction rounds while maintaining computational feasibility.
-For the base model, we used GPT-4o with instruction tuned agent, as it yielded the strongest overall performance.
+For the base model, we used GPT-4o with an instruction-tuned agent, as it yielded the strongest overall performance.
 3.5.4. Answering RQ4:
-In this research question (RQ), we conduct an ablation study by systematically removing each agent to assess its necessity.
+In this RQ, we conduct an ablation study by systematically removing each agent to assess its necessity.
 Since the security researcher agent is fundamental for identifying potential vulnerabilities, we retain it in all scenarios.
 Additionally, since the moderator primarily provides summaries rather than making explicit decisions on vulnerabilities, we do not include a scenario where the moderator is responsible for delivering the final verdict.
-In total, we evaluate four scenarios.
-First, the Full VulTrial (baseline) includes all four agents.
-Second, we exclude the moderator and use only the security researcher, code author, and review board.
-Third, we exclude the code author.
-Finally, we consider the effect of removing two agents at once.
-In this fourth scenario, we explore (a) only the security researcher and review board (i.e., no code author, no moderator), and (b) only the security researcher and code author (i.e., no moderator, no review board).
-For scenario 4(b), as the final decision is made by the code author agent, we classify the code as vulnerable if the code author mitigates at least one of the potential vulnerabilities.
+We evaluated four scenarios: (1) the full VulTrial (baseline) includes all four agents; (2) removing the moderator; (3) removing the code author; (4) removing two agents simultaneously, using either (a) only the security researcher and review board or (b) only the security researcher and code author.
+For scenario 4(b), as the final decision is made by the code author agent, we classify the code as vulnerable if it mitigates at least one of the potential vulnerabilities.
 By comparing performance across these scenarios, we can identify which agents play the most critical roles.
@@
 This ablation would highlight the necessity of each agent in VulTrial.
+3.6. Implementation Details
+We built VulTrial on top of AgentVerse (chen2024agentverse), a recently developed framework that simplifies the deployment of multiple LLM-based agents.
+For single-agent baselines, we used the prompt in PrimeVul (ding2024vulnerability), while for GPTLens (hu2023large), we employed the provided replication package (gptlens).
+Because the original GPTLens study focused on smart contracts, we adapted its prompt by replacing references to “smart contract” with “code” to generalize the use case.
+We used the best settings reported in their study, including two auditor agents and a top-$k$ value of 3.
+We counted the results as vulnerable if they got a final score of at least 5 (out of 9).
+All GPT-based models were accessed via the OpenAI API, which was also used to generate instruction tuning data.
+For the fine-tuning itself, we used OpenAI’s built-in dashboard feature for custom GPT-3.5 and GPT-4o instruction tuning (openai2024gpt4o_finetuning).
+To make the results more deterministic, we set the temperature to 0 for both our single-agent baselines and VulTrial.
+For GPTLens (hu2023large), we followed the same setting used in its evaluation (0.7 for the auditor agent and 0 for the critic agent).
+4. Results
+4.1. RQ1: VulTrial vs. Baselines
+Table 1 presents the performance of VulTrial against both single-agent (CodeBERT (feng2020codebert), CodeT5 (wang2021codet5), UniXCoder (guo2022unixcoder), LineVul (fu2022linevul), and Ding et al.’s CoT prompts (ding2024vulnerability)) and multi-agent (GPTLens (hu2023large)) baselines.
+In addition to the pair-wise metrics (ding2024vulnerability) (P-C, P-V, P-B, P-R, Error), we also report traditional metrics (P, R, FPR).
 Table 1.
-Results of various LLMs on PrimeVul paired functions.
+Results of various vulnerability detection methods on PrimeVul paired functions.
 An upward-pointing arrow ($\uparrow$) indicates that a higher score is better, while a downward-pointing arrow ($\downarrow$) indicates that a lower score is better.
-Model Agents Method P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
-Ding et al.’s CoT (Ding et al., 2025b) 6.21 4.79 83.51 5.5
-Single-Agent Finetuned 1.24 5.32 90.96 2.48
-GPTLens (Hu et al., 2023) 4.60 89.20 0.69 5.51
-GPT-3.5 Multi-Agent VulTrial 15.63 9.20 60.92 14.25
-Single-Agent Ding et al.’s CoT (Ding et al., 2025b) 9.20 9.89 74.26 6.67
-GPTLens (Hu et al., 2023) 10.11 55.40 28.05 6.44
-GPT-4o Multi-Agent VulTrial 18.62 41.15 28.74 11.49
-  • $\dagger$ Results taken from previous study by Ding et al. (Ding et al., 2025b).
-3.6. Implementation Details
-We built VulTrial on top of AgentVerse (Chen et al., 2024), a recently developed framework that simplifies the deployment of multiple LLM-based agents.
-For single-agent baselines, we used the prompt in PrimeVul (Ding et al., 2025b), while for GPTLens (Hu et al., 2023), we employed the provided replication package (Hu et al., 2022).
-Because the original GPTLens study focused on smart contracts, we adapted its prompt by replacing references to “smart contract” with “code” to generalize the use case.
-We used the best settings reported in their study, including two auditor agents, a top-$k$ value of 3, and the criteria for vulnerable with a final score of at least 5 (out of 9).
-All GPT-based models were accessed through the OpenAI API.
-For instruction tuning GPT-3.5 and GPT-4o, we utilized OpenAI’s official fine-tuning tools (OpenAI, 2024a).
-To make the results more deterministic, we set the temperature to 0 for both our single-agent baselines and VulTrial.
-For GPTLens (Hu et al., 2023), we followed the same setting used in its evaluation (0.7 for the auditor agent and 0 for the critic agent).
-4. Results
-4.1. RQ1: VulTrial vs. Single-Agent and Multi-Agent Baselines
-In this RQ, we evaluated VulTrial against single-agent (using Ding et al.
-CoT prompts (Ding et al., 2025b)) and multi-agent (GPTLens (Hu et al., 2023)) baselines across several LLMs.
-We utilized GPT-3.5 and GPT-4o for the base model in VulTrial, single-agent, and multi-agent baselines (see Section 3).
-Table 1 presents the results, with P-C indicating pairs classified correctly for both functions (higher is better).
-GPT-3.5.
-Among all models, GPT-3.5 saw the largest boost when applying VulTrial.
-Its P-C score increased by 151.69%, over the single-agent baseline.
-Likewise, P-B dropped from 83.51 to 60.92, which highlights the number of both codes in the pair that were predicted as benign are less in VulTrial.
-Furthermore, VulTrial also improves upon the performance of GPTLens by 239.78% on P-C.
-In most cases, GPTLens predicts functions as vulnerable, as shown P-V score (i.e., pairs incorrectly labeled as vulnerable) of 89.2, which is much higher compared to VulTrial (9.2).
-GPTLens is configured to output the top-3 potential vulnerabilities per function, following the best setup from the original study (Hu et al., 2023).
-Since auditor agent runs twice, this results in up to six potential vulnerabilities per function.
+Model Agents Method P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$ Error $\downarrow$ P $\uparrow$ R $\uparrow$ FPR $\downarrow$
+CodeBERT (feng2020codebert) Single-Agent Fine-tuned full (ding2024vulnerability) 5 37 386 7 430 0.49 0.10 0.10
+Fine-tuned pair 0 435 0 0 435 0.50 1.00 1.00
+CodeT5 (wang2021codet5) Single-Agent Fine-tuned full (ding2024vulnerability) 0 57 373 5 435 0.48 0.13 0.14
+Fine-tuned pair 0 434 1 0 435 0.50 0.99 0.99
+UniXCoder (guo2022unixcoder) Single-Agent Fine-tuned full (ding2024vulnerability) 4 26 403 2 431 0.52 0.07 0.06
+Fine-tuned pair 6 412 17 0 429 0.50 0.96 0.95
+LineVul (fu2022linevul) Single-Agent Fine-tuned§ (fu2022linevul) 7 89 336 3 428 0.51 0.22 0.21
+Fine-tuned pair 0 434 1 0 435 0.50 0.10 1.00
+Single-Agent Ding et al.’s CoT (ding2024vulnerability) 18 16 381 20 417 0.49 0.08 0.08
+GPTLens (hu2023large) 20 388 3 24 415 0.50 0.94 0.95
+GPT-3.5 Multi-Agent VulTrial 68 40 265 62 367 0.51 0.25 0.23
+Single-Agent Ding et al.’s CoT (ding2024vulnerability) 40 43 323 29 395 0.53 0.19 0.17
+GPTLens (hu2023large) 44 241 122 28 391 0.51 0.65 0.62
+GPT-4o Multi-Agent VulTrial 81 179 125 50 354 0.53 0.60 0.52
+  • $\dagger$ Fine-tuned on full PrimeVul train and val sets (4,862 vuln / 170,935 benign).
+    $\ddagger$ Fine-tuned on paired PrimeVul train and val sets (3,789 vuln / 3,789 benign).
+  • $\S$ Fine-tuned model from prior work (fu2022linevul).
+Deep learning-based Method.
+[…]
+These results highlight the limitations of traditional DL models in capturing subtle vulnerability patterns at the function-pair level.
+While DL-based models are more lightweight and computationally efficient, our results demonstrate that this efficiency comes at a cost in actionable performance.
+For example, compared to the highest performing DL baseline, LineVul, achieves only 7 for P-C, with a modest FPR of 0.21 and a recall of 0.22.
+VulTrial (using GPT-3.5) achieves a P-C of 68 (nearly 10× higher) while maintaining a comparable FPR and recall of 0.23 and 0.25.
+With GPT-4o, VulTrial pushes the P-C even further to 81 (12× higher than LineVul), with recall reaching 0.60.
+Although this comes at the cost of a higher FPR (0.52), the trade-off results in vastly more correct pairwise classifications and a significantly higher yield of true vulnerabilities.
+LLM-based Method (GPT-3.5).
+The single-agent CoT prompt obtains modest accuracy (P-C=18, FPR=0.08) but misses most vulnerabilities (R=0.08).
+Meanwhile, GPTLens boosts recall to 0.94; however, it comes with a high FPR of 0.95 with a small number of correct pairs (20).
+VulTrial obtains 68 correct pairs, which is 3.8x the CoT baseline and 3.4x GPTLens, while keeping FPR to 0.23.
+GPTLens is configured to output the top-3 potential vulnerabilities per function, following the best setup from the original study (hu2023large).
+Since the auditor agent runs twice, this results in up to six potential vulnerabilities per function.
 Indeed, most auditor outputs list six vulnerabilities, with an average of 5.6 and a median of 6.
-Meanwhile, VulTrial produce an average of 1.74 vulnerabilities per function, with a median of 2.
+Meanwhile, VulTrial produces an average of 1.74 vulnerabilities per function, with a median of 2.
 In other words, VulTrial produces fewer vulnerabilities compared to GPTLens, but it has more precise vulnerability detection, shown by a higher P-C.
-GPT-4o.
-Using GPT-4o as the base model, VulTrial also achieves the best performance.
-Specifically, it outperforms the single-agent baseline by 102.39% and the multi-agent baseline by 84.17% in the P-C metric.
-Across all methods (VulTrial, single-agent, and multi-agent), performance under GPT-4o is higher than with GPT-3.5.
-This aligns with OpenAI’s claim that GPT-4o offers higher performance than GPT-3.5, as well as with previous benchmarking studies that compare these models (OpenAI, 2024b; Wang et al., 2024; Günay-polatkan and Sığırlı, 2024; Kipp, 2024).
-For the single-agent baseline with GPT-4o, similar to its GPT-3.5 performance, it exhibited a high P-B score (pairs incorrectly classified as benign), suggesting a tendency to overlook vulnerabilities.
-However, in GPT-4o the P-B score dropped to 74.26, compared to GPT-3.5 (83.51).
-In contrast, VulTrial significantly reduced P-B to 28.74, a lower value than both the single-agent GPT-4o and its own performance under GPT-3.5 (60.92).
-This demonstrates VulTrial’s improved ability to correctly predict benign code with GPT-4o.
-Examining P-R (pairs where the predicted classification is reversed from the ground truth), VulTrial with GPT-4o also showed an improvement, reducing P-R by 24.02% compared to its performance with GPT-3.5.
+LLM-based Method: GPT-4o.
+VulTrial achieves the highest number of pair-wise correct predictions (P-C=81), outperforming both the CoT baseline (P-C=40) and GPTLens (P-C=44).
+This is 2.03× the performance of the single-agent baseline and 1.84× that of GPTLens.
+These gains highlight VulTrial’s effectiveness at correctly identifying both the vulnerable and benign functions within a pair that have subtle differences.
+In terms of traditional classification metrics, VulTrial achieves a precision of 0.53, on par with the CoT baseline (0.53) and slightly above GPTLens (0.51).
+Its recall is 0.60, lower than GPTLens (0.65), but significantly higher than CoT (0.19).
+The FPR for VulTrial is 0.52, which is lower than GPTLens (0.62) but higher than CoT (0.17).
+This trade-off illustrates VulTrial’s balanced behavior; it identifies more true vulnerabilities than CoT while maintaining a lower FPR than GPTLens.
+In terms of P-B, which measures how often both functions in the pair were incorrectly classified as benign, VulTrial (P-B=125) is significantly lower than the CoT baseline (P-B=323) and comparable to GPTLens (P-B=122).
+Despite these improvements, VulTrial does not outperform the alternatives on every metric.
+GPTLens still achieves the highest recall, making it preferable in scenarios where catching as many vulnerabilities as possible is critical.
 In the GPTLens experiment using GPT-4o, the first agent highlights a similar number of potential vulnerabilities as it did under GPT-3.5, with a median of 6 and a mean of 5.76.
 Meanwhile, VulTrial with GPT-4o reports fewer vulnerabilities (median 4, mean 4.11) compared to GPTLens.
-Despite highlighting fewer vulnerabilities, VulTrial achieved higher P-C and lower P-B scores than GPTLens, indicating overall better performance.
+Despite highlighting fewer vulnerabilities, VulTrial achieved higher P-C and lower P-B scores than GPTLens.
 This suggests that the additional agents in VulTrial, can effectively distinguish between justified and unjustified vulnerability claims, leading to more accurate final predictions.
-In Fig.
-2, we show an example conversation from VulTrial.
-The code snippet under discussion is flagged for failing to validate negative indices, which allow an attacker to read arbitrary data by inputting out-of-range values in indices.
-The security researcher agent identifies two vulnerabilities: (1) Out-of-Bounds Read (CWE-125), which is valid, and (2) Type Handling, which is misleading since the existing code already handles unsupported types safely.
-After reviewing these findings, the code author agent agrees with the Out-of-Bounds Read assessment but refutes the Type Handling concern, which shows the agent can distinguish confirmed risks from exaggerated claims.
-The moderator agent then summarizes key points from both agents, which aim to help review board agent in making the decision.
-Based on these information, the review board agent concludes that the Out-of-Bounds Read is a genuine vulnerability.
+Our qualitative analysis further illustrates how VulTrial’s structured, role-specific dialogue improves both fidelity and precision of vulnerability detection.
+Each agent in VulTrial has a specific role.
+The security researcher agent analyzes the function to identify vulnerabilities.
+For example, in Fig.
+2, the security researcher highlights two vulnerabilities: (1) an Out-of-Bounds Read due to unvalidated negative indices (CWE-125), and (2) a Type Handling issue.
+The first is a valid case, while the second reflects an overcautious concern since the existing code already handles unsupported types safely.
+After reviewing these findings, the code author agent agrees with the Out-of-Bounds Read assessment but refutes the Type Handling concern.
+The moderator agent then summarizes key points from both agents, which aim to help the review board agent in making the decision.
+Based on this information, the review board agent concludes that the Out-of-Bounds Read is a genuine vulnerability.
 Meanwhile, for the Type Handling issue, the review board agrees with the code author’s argument that the existing implementation is safe.
-However, they worried about potential future changes that could introduce vulnerabilities as highlighted by the security researcher agent.
-As a result, the board classifies this issue as partially valid with a recommendation to monitor.
-Answer to RQ1 Using GPT-3.5 as base LLMs, VulTrial improves P-C of single-agent and multi-agent baseline by 151.69% and 239.78%, respectively.
-Meanwhile, using GPT-4o as the base LLMs, VulTrial improve the baselines by 102.39% and 84.17%.
-4.2. RQ2: Instruction Tuning on Role-Specific Agents
-From the RQ1’s results, we observed that VulTrial using GPT-4o achieve the best performance.
+However, it’s concerned about potential future changes that could introduce vulnerabilities, as highlighted by the security researcher agent.
+As a result, this issue is classified as partially valid, which is considered benign since it falls below the threshold.
+This example illustrates how the agents collectively separate confirmed vulnerabilities from exaggerated ones, preventing over-reporting and guiding appropriate remediation.
+Another example of running VulTrial on vulnerable code can be found in Appendix H (ourreplicationpackageall).
+The security researcher agent flags the lack of strict checks; the code author proposes callbacks to mitigate the issue, and the review board ultimately rules it a high-severity, valid vulnerability requiring immediate fix.
+Here, both the security researcher and code author align, reinforcing the confidence of the final judgment.
+This demonstrates VulTrial’s capacity to not only support adversarial debate but also consensus when warranted.
+Conversely, in the example of the benign function that can be found in Appendix I (ourreplicationpackageall), the security researcher agent initially raises a concern about certificate validation.
+The code author agent points out that X.509 verification is already used, which the review board agent independently confirms.
+The issue is ruled invalid.
+This case highlights VulTrial’s strength in filtering false positives as the security researcher agent’s overstatement is corrected by technical context from the code author’s agent.
+In another example of benign function (Appendix J (ourreplicationpackageall)), the security researcher agent flags potential leakage via verbose logging.
+The code author proposes mitigations: increasing verbosity thresholds, securing log storage, and reviewing logged content.
+The moderator captures this exchange.
+After reviewing the evidence, the review board classifies the issue as partially valid with low severity (which is below the vulnerability threshold), opting to monitor rather than patch, thereby filtering out an overcautious report.
+Our qualitative analysis shows that the four agents contribute different kinds of knowledge: pattern detection (security researcher), code-context refutation or mitigation (code author), factual summarisation (moderator), and balanced final judgment (review board).
+This minimizes errors on both ends: they surface genuine threats while filtering out already-handled or overstated concerns, reducing false negatives and false positives.
+Answer to RQ1: VulTrial achieves the highest number of correct function pair predictions while maintaining a balanced precision and recall score.
+It more than doubles the number of fully‑correct function pairs relative to baselines, while avoiding the extreme false‑positive rates and the low recall.
+4.2. RQ2: Instruction Tuning on Agents
+From the RQ1’s results, we observed that VulTrial using GPT-4o achieves the best performance.
 In this RQ, we investigate whether instruction tuning on role-specific agents can further improve VulTrial’s overall performance.
-The results of this instruction tuning phase are shown in Table 2.
-Table 2. Results of VulTrial using GPT-4o with instruction tuned agents. SR represents the security researcher agent, CA the code author agent, M the moderator agent, and RB the review board agent.
-Instruction Tuning P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
-None 18.62 41.15 28.74 11.49
-Security researcher (SR) 12.41 35.17 42.99 9.43
-Code author (CA) 11.95 8.97 72.18 6.90
-Moderator (M) 22.07 35.40 30.80 11.72
-Review board (RB) 16.55 30.57 42.07 10.80
-SR + CA 7.82 9.20 76.09 6.90
-SR + CA + M 7.36 8.51 77.70 6.44
-SR + CA + MR + RB 10.34 10.57 70.11 8.97
+The results are shown in Table 2.
 GPT-4o Instruction Tuning.
-All the single instruction tuned variants of GPT-4o in VulTrial, covering the security researcher agent, code author agent, moderator agent, and review board agent, outperformed both the single-agent and multi-agent baselines.
+All the single instruction-tuned variants of GPT-4o in VulTrial, covering the security researcher agent, code author agent, moderator agent, and review board agent, outperformed both the single-agent and multi-agent baselines.
 Among these, tuning the moderator agent delivered the most substantial performance gain.
-Specifically, the moderator-tuned version increases the performance of the non-tuned version in VulTrial by 18.53% (P-C).
-Compared to the single-agent baseline, this instruction tuned version of VulTrial increases P-C by 139.89%.
-While for the multi-agent baseline, the improvement of the instruction tuned version of VulTrial increases P-C by 118.30%.
+Specifically, the moderator-tuned version increases the performance of the non-tuned version in VulTrial by 15 P-C scores.
+Compared to the single-agent baseline, this instruction-tuned version of VulTrial increases P-C score by 56.
+While for the multi-agent baseline, the improvement of the instruction-tuned version of VulTrial increases P-C score by 52.
+Table 2. Results of VulTrial using GPT-4o with instruction-tuned agents: SR (security researcher), CA (code author), M (moderator), and RB (review board).
+Instruction-tuned P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
+None 81 179 125 50
+Security researcher (SR) 54 153 187 41
+Code author (CA) 52 39 314 30
+Moderator (M) 96 154 134 51
+Review board (RB) 72 133 183 47
+SR + CA 34 40 331 30
+SR + CA + M 32 37 338 28
+SR + CA + M + RB 45 46 305 39
 Effect of Agent-Specific Tuning.
-Upon further analysis of the instruction tuned moderator agent, we observed that its summaries are more concise compared to the non-tuned version.
+Upon further analysis of the instruction-tuned moderator agent, we observed that its summaries are more concise compared to the non-tuned version.
 Specifically, the tuned version of the moderator agent provides a shorter summary than the non-tuned version, reducing the summary length for security researchers by 16.03% and for code authors by 16.58%.
@@
 Meanwhile, when the tuned VulTrial correctly classified code as vulnerable but the non-tuned version did not, the code author’s summary was 19.03% shorter.
-These observations suggest that the instruction tuned moderator agent can effectively distill key points from both the security researcher and code author, which help the review board agent to make more accurate final verdicts.
+These observations suggest that the instruction-tuned moderator agent can effectively distill key points from both the security researcher and code author, which help the review board agent to make more accurate final verdicts.
 We next examined the impact of instruction tuning on the security researcher and code author agents.
@@
 Similarly, instruction tuning the code author agent altered its mitigation-to-refutation ratio.
-While the mitigation to refutation ratio of non-tuned version was 2:1, the tuned version exhibited a 1:4 ratio.
-In other words, the tuned code author refuted vulnerabilities far more frequently.
+While the mitigation to refutation ratio of the non-tuned version was 2:1, the tuned version exhibited a 1:4 ratio.
+In other words, the tuned code author refuted vulnerabilities more frequently.
 From its tuning data, it learned to doubt potential vulnerabilities highlighted by security researcher.
-Although this approach can restrict over-diagnosing vulnerabilities, it also risks undervaluing genuine security concerns, as highlighted by an increase in P-B (cases where both snippets are incorrectly labeled benign).
-Despite outperforming single-agent baseline, this shows the delicate balance that has to be achieved by instruction tuning; While instruction tuning can sharpen an individual agent’s skills, it may introduce biases or extreme behaviors that undermine collective performance.
+Although this approach can restrict over-diagnosing vulnerabilities, it risks undervaluing genuine security concerns, as highlighted by an increase in P-B.
+Despite outperforming the single-agent baseline, this shows the delicate balance required for instruction tuning.
+While instruction tuning can sharpen an individual agent’s skills, it may introduce biases or extreme behaviors that undermine collective performance.
+Table 3. Results of VulTrial using GPT-4o across different numbers of discussion rounds.
+Model #Round P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
+1 96 154 134 51
+2 87 173 122 53
+GPT-4o 3 81 194 110 50
+  • $\ddagger$ Instruction tuning for moderator agent.
 Combining Agent-Specific Tuning.
@@
 These combinations did not surpass the performance achieved by tuning only the moderator agent.
-For instance, when all agents were instruction tuned, the average number of identified vulnerabilities dropped from 4.11 to 2.60, and the refutation-to-mitigation ratio shifted to 1:4.
-This imbalance corresponded to a higher P-B rate of 70.11.
-Answer to RQ2 Instruction tuning for moderator agent using a small dataset (50 paired samples) yields the most significant improvements.
-It enhances performance by 18.53% compared to the non-tuned version and increases P-C by 139.89% and 118.30% over the single-agent and multi-agent baselines, respectively.
-4.3. RQ3: Effect of Multiple Rounds of Discussion
+For instance, when all agents were instruction-tuned, the average number of identified vulnerabilities dropped from 4.11 to 2.60, and the refutation-to-mitigation ratio shifted to 1:4.
+This imbalance corresponded to a higher P-B (331).
+These lower results may stem from the instruction tuning setup, where the review board is trained on histories that include inputs from other agents.
+In the tuning data, the review board often learns to “follow” the code author’s position (especially strong arguments), which is aligned with the ground truth.
+Once the code author is instruction-tuned and becomes more assertive in refuting vulnerabilities, the review board (also fine-tuned) tends to side with the author.
+Future work should decouple agent training, for example, including cases where the review board correctly identifies a vulnerability despite strong refutations, to reduce over-alignment and improve decisions.
+Answer to RQ2: Instruction tuning for moderator agent using 50 paired samples yields the most significant improvements.
+It improves P-C by 15, 56, and 52 compared to the non-tuned version, single-agent, and multi-agent baselines, respectively.
+4.3. RQ3: Effect of Multi Rounds Discussion
 In the previous RQs, VulTrial used a single round of interaction among the agents.
-In this RQ, we investigate whether additional discussion rounds could enhance performance.
-The results of these multiple rounds experiments are shown in Table 3.
-The best result is achieved by using a single round interaction.
-Table 3. Results of VulTrial using GPT-4o across different numbers of discussion rounds.
-Model #Round P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
-1 22.07 35.40 30.80 11.72
-2 20.00 39.77 28.05 12.18
-GPT-4o 3 18.62 44.60 25.29 11.49
-  • $\ddagger$ Instruction tuning for moderator agent.
+In this RQ, we investigate whether additional discussion rounds enhance performance.
+Table 3 shows the results of these multiple rounds experiments, showing that the best result is achieved using a single round interaction.
 For GPT-4o, increasing the number of rounds actually led to a decrease in overall performance.
-Specifically, the proportion of pairs incorrectly categorized as vulnerable (P-V) increases with more rounds.
+Specifically, P-V increases with more rounds.
 From further analysis, we found that the security researcher agent suggested fewer potential vulnerabilities in later rounds, with an average of 4.22 per code in Round 1, decreasing to 3.55 in Round 2 and 3.43 in Round 3.
 This decline is also driven by the code author agent’s refutations.
-In 30.88% out of the total prediction, the security researcher removed a previously identified vulnerability after the code author disputed it.
+In 30.88% of the total predictions, the security researcher removed a previously identified vulnerability after the code author disputed it.
 Meanwhile, the code author agent became progressively more inclined to acknowledge vulnerabilities, with its mitigation-to-refutation ratio shifting from 1.73:1 in Round 1, to 2.5:1 in Round 2, and finally 3:1 in Round 3.
 Although the review board agent did not blindly accept arguments from either side (for instance, during the three-round setup, in 100 out of 868 code instances, review board agent overruled the code author’s refutation and decided the vulnerability to be valid), the overall imbalance introduced by these interactions still raised P-V and lower P-C rates.
-Answer to RQ3 Using GPT-4o in VulTrial outperforms both single-agent and multi-agent baselines across different rounds of agent interaction (i.e., 1, 2, and 3).
+Answer to RQ3: Using GPT-4o in VulTrial outperforms both single-agent and multi-agent baselines across different rounds of agent interaction (i.e., 1, 2, and 3).
 However, performance declines when extending beyond a single discussion round.
-Table 4. Results of ablation study on agent in VulTrial using GPT-4o with moderator agent tuned.
-Agents P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
-Base 22.07 35.40 30.80 11.72
-No moderator 7.36 8.51 77.70 6.44
-No code author 14.48 63.68 11.95 9.89
-No moderator+code author 14.48 61.84 14.48 9.20
-No moderator+review boards 2.07 93.56 1.15 3.22
 4.4. RQ4: Ablation Study
@@
 Overall, we find that all four agents are essential for achieving the best performance.
-Removing the moderator agent from VulTrial reduces P-C from 22.07 to 7.36.
+Removing the moderator agent from VulTrial reduces P-C from 96 to 32.
 This result highlights the moderator’s critical role in summarizing arguments to help review board agent in making the decision.
 This finding is also consistent with the results of RQ2 (see Section 4.2), where instruction tuning the moderator yielded the largest performance gain.
-Removing both the moderator and review board agent causes P-C to further drop 2.07, while P-V (pairs both labeled vulnerable) increase to 93.56.
+Removing both the moderator and review board agent causes P-C to further drop to 9, while P-V (pairs both labeled vulnerable) increases to 407.
 Highlighting that most of the codes are predicted as vulnerable by the code author.
@@
 Thus, the review board is essential for ensuring balanced final verdicts.
-Removing the code author agent from VulTrial reduces P-C to 14.48 and increases P-V to 63.68.
+Removing the code author agent from VulTrial reduces P-C to 63 and increases P-V to 277.
 Without the code author’s counterarguments, the review board agent has a limited perspective and frequently accepts vulnerabilities identified by the security researcher agent.
-Removing both the moderator and code author produces similar results (P-C: 14.48, P-V: 61.84).
+Removing both the moderator and code author produces similar results (P-C=63, P-V=269).
 This underscores the crucial role of the code author in ensuring balanced judgments while also demonstrating that the moderator agent becomes redundant without opposing arguments to mediate.
-Answer to RQ4 Our findings confirm that every agent in VulTrial is necessary to achieve the best performance.
+Answer to RQ4: Our findings confirm that every agent in VulTrial is necessary to achieve the best performance.
+Table 4. Results of ablation study on agent in VulTrial using GPT-4o with moderator agent tuned.
+Agents P-C $\uparrow$ P-V $\downarrow$ P-B $\downarrow$ P-R $\downarrow$
+Base 96 154 134 51
+No moderator 32 37 338 28
+No code author 63 277 52 43
+No moderator+code author 63 269 63 40
+No moderator+review boards 9 407 5 14
 5. Discussion
 Cost vs. Performance Trade-offs
-By assigning specialized roles to agents, mirroring real-world scenarios, VulTrial can facilitate deeper discussions by incorporating diverse rationales from different parties, even when using small-scale LLMs.
+By assigning specialized roles to agents, mirroring real-world scenarios, VulTrial can facilitate deeper discussions by incorporating diverse rationales from different parties.
 However, this multi-agent approach inherently incurs additional costs due to the increased number of input and output tokens.
-Interestingly, while GPT-4o is generally more expensive and performs better than GPT-3.5 as a single-agent, with respective P-C 9.2 and 6.21, using a multi-agent approach (VulTrial) with GPT-3.5 outperforms GPT-4o’s single-agent performance by 69.89%.
+Interestingly, while GPT-4o is generally more expensive and performs better than GPT-3.5 as a single-agent, using VulTrial with GPT-3.5 outperforms GPT-4o’s single-agent P-C by 28.
 This approach also reduces costs, requiring only \$7.46 compared to GPT-4o’s single-agent cost of \$8.10 for the whole testing data.
@@
 Therefore, careful consideration is important when selecting models to optimize both cost efficiency and detection performance.
-Multi-Agents Performance
-In both experiments using GPT-3.5 and GPT-4o as base models for GPTLens (see Section 4.1), we observed that a large portion of the codes were predicted as vulnerable, as shown by high P-V scores (incorrectly predicted as vulnerable) of 89.2 and 41.15, respectively.
-This behavior may be due to the critic agent (final agent) assigning high scores on potential vulnerabilities flagged by the auditor agent (first agent).
-To get the final prediction, GPTLens also utilizes a threshold to determine whether a function is vulnerable.
-The final score, ranging from 0 to 9, is computed as $0.5\times\text{correctness}+0.25\times\text{severity}+0.25\times\text{profitability}$, with correctness, severity, and profitability taken from the critic agent output.
-The final scores below 5 are considered benign.
-When GPTLens uses GPT-3.5, the median final score is 7.25 with mean 7.08, indicating that most code is labeled as vulnerable.
-Similarly, when GPTLens uses GPT-4o, the median final score is 6.75 with mean 6.67, which means the majority of the code prediction also exceeds the benign threshold.
-Meanwhile, under GPT-4o, VulTrial demonstrated a more balanced results, as indicated by a mitigation-to-refutation ratio of 2:1 in the code author agent’s responses.
-Furthermore, we found that in 101 benign and 253 vulnerable codes, the code author agent refuted at least one potential vulnerability initially flagged by the security researcher (i.e., disagreement between code author and security researcher), yet the final label remained correct.
-These cases, accounting for 40.78% of the total codes, demonstrate VulTrial’s effectiveness in resolving disagreements.
-Enhancing Role-Specific Agent
-Tuning the moderator agent in VulTrial using GPT-4o enhances single-agent performance by 139.89% and improves the non-tuned version of VulTrial by 18.52%.
-This underscores the importance of instruction tuning for role-specific agents.
-While instruction tuning on all four agents in VulTrial would cost four times more than tuning a single-agent.
-In this study, we used only a sample of 50 pairs for instruction tuning, which kept the cost at approximately \$17 for each agent.
-This is significantly lower than using the instruction tuning setting done in PrimeVul (Ding et al., 2025b), which would cost approximately more than \$500 for GPT-4o (¿15,000 samples).
+Enhancing Role-Specific Agents
+Tuning the moderator agent in VulTrial using GPT-4o improves the baselines and non-tuned version of VulTrial.
+These results highlight the potential for improving performance by tuning role-specific agents.
+Our study used only a sample of 50 pairs for instruction tuning, keeping the cost at approximately \$17 for each agent.
+This is significantly lower than using the instruction tuning setting done in PrimeVul (ding2024vulnerability), which would cost more than \$500 for GPT-4o (¿15,000 samples).
+Analysis on Instruction Tuning Data.
+Comparing the 50 fine-tuning pair samples from the validation dataset with the 96 successful VulTrial pairs, we found that 39 projects from the successful cases do not appear in the fine-tuning set.
+Only 8 projects are shared.
+Of these, 3 projects contain only commits that occurred after the corresponding training commits.
+In total, only 11 successful pairs involve commits made earlier than the latest commit used in training.
+The fine-tuning data includes 24 unique CWEs, 20 of which also appear in the successful set.
+12 CWEs in the successful set are not covered by the fine-tuning data.
+These results show minimal overlap in terms of projects, commit timing, and CWE types, suggesting that the model’s performance is not a result of memorization.
+Table 5. Results of running VulTrial in the wild.
+Project Funcs analyzed Alarms raised False positives True positives
+Superset (superset) 47 0 0 0
+DataHub (datahub) 160 2 0 2
+InvenTree (inventree) 11 1 0 1
+VulTrial in the Wild
+To assess VulTrial in real-world settings, we conducted a pilot experiment on selected open-source projects.
+Specifically, we targeted well-maintained repositories with active security practices.
+To identify such projects, we filtered GitHub repositories that include a SECURITY.md file and have over 1,000 stars.
+We chose the projects that have a SECURITY.md file, as it typically indicates that maintainers are attentive to security concerns and have set up a process for reporting vulnerabilities.
+This selection process yielded 196 repositories.
+From this set, we randomly chose three projects for analysis.
+For each project, we analyzed recent commits, extracted modified functions, and applied VulTrial with its default rule set, no project-specific tuning.
+The results of this experiment are summarized in Table 5.
+We manually inspected the flagged vulnerabilities, attempted exploit generation, and reported confirmed issues to the respective maintainers if the vulnerable code was used in the project (i.e., not dead code).
+In InvenTree, we discovered a denial-of-service vulnerability in the built-in label_sheet plugin.
+The plugin accepts a skip parameter with no upper bound, allowing an attacker to allocate excessively large Python lists and exhaust server memory.
+This issue affects any authenticated user who can request labels and has been assigned CVE-2025-49000.
+In DataHub, VulTrial identified two functions with resource-exhaustion risks (CWE-400).
+One function is still actively used, while the other is no longer in use but remains in the codebase.
+We disclosed the first issues following instructions on their SECURITY.md, and the DataHub developers confirmed the findings.
+These results demonstrate that VulTrial is capable of detecting previously unknown vulnerabilities in the wild.
+To reduce the burden on maintainers, we adopted a conservative reporting strategy and only reported issues we could confidently confirm as true vulnerabilities.
+We have made the results, vulnerability reports, and exploits available (ourreplicationpackage).
+Generalizability of VulTrial
+In our main experiments, VulTrial, using OpenAI’s models, outperforms the baseline methods.
+Next, to investigate its generalizability, we further evaluated it with a different model family, specifically, LLaMA-3.1-8B (dubey2024llama).
+Table 6 shows the results, indicating the consistent performance of VulTrial even with a different model family.
+CoT struggles under LLaMA-3.1-8B, with only 4 correct predictions (P-C) and a recall of 0.02, largely due to a high number of benign misclassifications (P-B = 425).
+VulTrial improves recall to 0.48 and raises P-C to 89, although it has an increased false-positive rate (FPR = 0.45).
+This suggests that VulTrial favors broader vulnerability coverage, potentially at the cost of mislabeling some benign pairs.
+When compared with GPTLens, VulTrial presents a different trade-off.
+GPTLens achieves higher recall (0.55) and fewer benign misclassifications (P-B = 103), but it also produces more P-V (158) and a higher FPR (0.58).
+VulTrial, in contrast, has more correct predictions overall (P-C = 89 vs.
+80) and higher precision (0.52 vs.
+0.49), while maintaining a lower false-positive rate (0.45).
+These differences indicate that VulTrial may provide a more conservative alternative to GPTLens, with improved balance across evaluation metrics.
+Overall, these findings suggest that VulTrial’s multi-agent reasoning approach can generalize to other models.
+Table 6. Results of VulTrial, single-agent, and multi-agent baselines with LLaMA-3.1-8B as the base model.
+[…]
+CoT(ding2024vulnerability) 4 3 425 3 431 0.54 0.02 0.01
+GPTLens(gptlens) 80 158 103 94 355 0.49 0.55 0.58
+VulTrial 89 120 150 76 346 0.52 0.48 0.45
+Evaluating Explanation
+VulTrial provides explanations for each agent’s decisions, helping developers understand why a function is classified as vulnerable or benign rather than receiving a simple binary outcome.
+Because multiple agents contribute, developers may need to read the full multi-agent discussion to grasp the reasoning.
+To improve readability, we use GPT-4o to synthesize all agent reasoning into a single coherent explanation.
+We compare the synthesized VulTrial explanations with CoT-based ones using four metrics (i.e., completeness, informativeness, clarity, and actionability), each scored on a 1-5 scale (widyasari2025explaining; degen2022explain; zhou2021evaluating; velmurugan2025developing).
+For fairness, we evaluate only cases where the P-C intersections of VulTrial and CoT match under GPT-4o.
+We conduct automated evaluation using a closed-source LLM from a different family (Claude-Haiku-4.5 (anthropic2025claudeHaiku)) and manual evaluation by three external annotators on 20 instances, without revealing explanation sources.
+Both evaluations show that synthesized VulTrial explanations outperform CoT across all metrics (Table 7).
+All generated explanations and evaluation results are available in the replication package (ourreplicationpackageall).
 6. Threats to Validity
 External Validity.
-An external threat to our study is the generalizability of our findings.
-In this study, we focused the evaluation using PrimeVul pair data, which is a challenging recently-released vulnerability detection task on C/C++ programming language.
+In this study, we focused the evaluation using PrimeVul pair data, which is a challenging recently-released vulnerability detection task on C/C++.
 We did not use other datasets as they may not be free from bias that have been uncovered by Ding et al.
-(Ding et al., 2025b).
+(ding2024vulnerability).
 In the future, we plan to investigate VulTrial performance using other clean datasets.
+Another threat is the risk of data leakage, as ChatGPT’s training data is not publicly disclosed.
+We mitigate this by evaluating VulTrial using LLaMA-3.1-8B, whose training cutoff (Dec 2023) is before PrimeVul’s GitHub release (Mar 2024).
+VulTrial also identifies previously unknown real-world vulnerabilities, including CVE-2025-49000, showing it can generalize beyond training data.
+Therefore, we believe the threats are minimal.
 Construct Validity.
-Threats to construct validity are from the choice of metrics in our study.
-In this study, we utilize P-C, P-B, P-V, P-R (Section 3.4).
-These metrics are designed for a pair-wise evaluation introduced in PrimeVul (Ding et al., 2025b).
-In a textually similar pair of functions, the metrics assesses an approach’s ability to correctly identify both the presence and absence of vulnerabilities.
-Consequently, we believe that the threat to construct validity is minimal.
+In this study, we use both pair-wise metrics and traditional metrics in our experiments (Section 3.4).
+These metrics provide a reliable measure of comparative performance, minimizing the threat to construct validity.
+In our experiments, VulTrial uses a single underlying model for all agents (e.g., GPT-4o, and LLaMA-3.1-8B), we rely on prompt-based role specialization to induce diverse behaviors (e.g., prosecutor vs.
+defense attorney).
+Prior works (wei2022chain; chatterjee2024posix; hu2023large; shinn2023reflexion; wu2024autogen) have shown that such prompt-induced roles can lead to qualitatively different reasoning paths even within the same model.
+In future work, we plan to use different LLMs for each agent role to further enhance diversity.
 Internal Validity.
@@
 To mitigate the risk, we have carefully checked the experiments that we run in our study.
-The code, prompt, and results are also included in our replication package: https://figshare.com/s/1514bc9a7aa64b46d94e.
-Furthermore, to reduce the randomness effect of the experiment, we set the temperature to 0 to make the results more deterministic.
+The code, prompt, and results are also included in our replication package (ourreplicationpackageall).
+Furthermore, we set the temperature to 0 to reduce the randomness effect of the experiment.
 We also did the preliminary evaluation where we ran the VulTrial on testing data three times and found that the standard deviation is low (0.7).
+Table 7. Results of automatic and manual evaluations of the generated explanation from CoT and synthesized VulTrial.
+CoT / Synthesized VulTrial
+Completeness Clarity Actionability Inform.
+Claude (anthropic2025claudeHaiku) 3.15 / 4.10 3.60 / 3.85 2.15 / 4.00 2.80 / 3.95
+Human 3.83 / 4.32 3.72 / 4.20 2.37 / 4.52 3.12 / 3.63
 7. Related Work
 Single-Agent Vulnerability Detection.
-A broad range of studies have explored prompting and CoT for vulnerability detection (Tamberg and Bahsi, 2024; Ullah et al., 2023; Zhou et al., 2024; Nong et al., 2024; Zhang et al., 2024; Ding et al., 2025b; Steenhoek et al., 2024).
-Tamberg and Bahsi (Tamberg and Bahsi, 2024) benchmarked LLMs using advanced prompting techniques, including CoT, tree-of-thought, and self-consistency, finding that CoT with GPT-4 achieved the best performance and outperformed static analysis tools.
+Recent work has explored both DL-based and LLM-based methods for vulnerability detection.
+DL approaches, such as LineVul (fu2022linevul), achieve high recall with the use of transformer-based line-level classification.
+Ni et al.
+(ni2023distinguishing) proposed a semantic representation learning method, while Steenhoek et al.
+(steenhoek2024dataflow) incorporated dataflow.
+In contrast, LLM-based methods have gained traction for their strong results (gptlens; zhou2024large_emerging).
+Other studies have explored prompting and CoT (tamberg2024harnessing; ullah2023llms; zhou2024large_emerging; nong2024chain; zhang2024prompt; ding2024vulnerability; steenhoek2024comprehensive).
+Tamberg and Bahsi (tamberg2024harnessing) benchmarked LLMs using advanced prompting techniques, including CoT, tree-of-thought, and self-consistency, finding that CoT with GPT-4 achieved the best performance and outperformed static analysis tools.
 Similarly, Zhou et al.
-(Zhou et al., 2024) demonstrated that GPT-4 with prompting outperforms fine-tuned CodeBERT.
+(zhou2024large_emerging) demonstrated that GPT-4 with prompting outperforms fine-tuned CodeBERT.
 Ullah et al.
-(Ullah et al., 2023) evaluated commercial LLMs across various prompting strategies, including CoT, and found promising detection performance.
-However, they noted often unfaithful reasoning in LLMs leads to incorrect classifications.
-Recent studies have further highlighted the limitations of CoT (Nong et al., 2024; Zhang et al., 2024; Steenhoek et al., 2024).
+(ullah2023llms) evaluated commercial LLMs across various prompting strategies, including CoT, and found promising detection performance.
+Recent studies have further highlighted the limitations of CoT (nong2024chain; zhang2024prompt; steenhoek2024comprehensive).
 Notably, PrimeVul’s seminal work found that GPT-3.5 and GPT-4 perform poorly on PrimeVul’s pair test set, reinforcing concerns about CoT’s reliability in vulnerability detection.
-In this work, we use CoT as a baseline and show that our approach, VulTrial, significantly outperforms it across various LLMs on PrimeVul.
+Efforts to enhance LLM-based methods include finetuning and retrieval-augmented generation (RAG).
+Shestov et al.
+(shestov2025finetuning) showed that finetuning boosts recall but is costly and dataset-sensitive.
+Du et al.
+(du2024vul) proposed Vul-RAG, which improves factual grounding via structured knowledge injection.
+Sun et al.
+(sun2024llm4vuln) introduced LLM4Vuln, a framework that evaluates and decouples LLM reasoning capabilities.
+In this work, we use DL-based methods and CoT as baselines and show that VulTrial outperforms them on PrimeVul.
 Multi-Agent Vulnerability Detection.
-Multi-agent frameworks can enhance LLM-based vulnerability detection by enabling collaboration and specialization among agents (He et al., 2024).
-For example, EvalSVA (Wen et al., 2024) introduces a multi-agent framework where agents act as security experts that improves over single-agent LLMs.
-Similarly, MuCoLD (Mao et al., 2024) adopts a multi-agent setup where agents assume roles such as tester and developer, which collaboratively analyze code.
+Multi-agent frameworks can enhance LLM-based vulnerability detection by enabling collaboration and specialization among agents (he2024llm).
+EvalSVA (wen2024evalsva) introduces a multi-agent framework where agents act as security experts.
+Similarly, MuCoLD (mao2024multi) agents assume roles such as tester and developer, which collaboratively analyze code.
 In contrast, VulTrial employs a structured multi-agent design inspired by a courtroom, where agents take on distinct roles such as prosecutor, defender, and judge.
-This courtroom analogy encourages adversarial debate and balanced reasoning that leads to robust decision-making.
-We excluded these recent works as baselines because the authors did not provide a replication package for reusing their agents, and the prompt information in the paper was insufficient to replicate the study.
-Other multi-agent-based approaches have been explored for smart contract auditing (Wei et al., 2024; Ma et al., 2024; Hu et al., 2023).
-LLM-SmartAudit (Wei et al., 2024) uses collaborative decision-making with role-play and reasoning to detect vulnerabilities in smart contracts.
-Similarly, iAudit (Ma et al., 2024) assigns specific roles, such as detector, reasoner, and critic, to guide the analysis of Solidity smart contracts.
-GPTLens (Hu et al., 2023) employs multiple LLM-based agents acting as independent prosecutors to identify vulnerabilities, with a critic agent reviewing and ranking the findings.
+This encourages adversarial debate and balanced reasoning that leads to robust decision-making.
+We excluded these recent works as baselines due to the lack of replication packages and insufficient prompt details to replicate the study.
+Other multi-agent-based approaches have been explored for smart contract auditing (wei2024llm; ma2024combining; hu2023large).
+LLM-SmartAudit (wei2024llm) uses collaborative decision-making with role-play and reasoning to detect vulnerabilities in smart contracts.
+Similarly, iAudit (ma2024combining) assigns specific roles, such as detector, reasoner, and critic, to guide the analysis of Solidity smart contracts.
+GPTLens (hu2023large) employs multiple LLM-based agents acting as independent prosecutors to identify vulnerabilities, with a critic agent reviewing and ranking the findings.
 Our work differs from these studies in two key ways.
@@
 Second, VulTrial adopts a structured, courtroom-inspired framework specifically tailored for general-purpose languages beyond smart contracts.
-As these studies focus on smart contracts, their scope differ from our work for general-purpose programming languages.
+As these studies focus on smart contracts, their scope differs from our work for general-purpose programming languages.
 Nonetheless, we found that GPTLens can be adapted to general-purpose vulnerability detection and included it as a baseline in our experiments (Section 3.3).
 Our results demonstrate that VulTrial substantially outperforms GPTLens and single-agent baselines on PrimeVul.
-8. Conclusion and Future Direction
-In this study, we proposed VulTrial, a novel multi-agent LLM-based approach for vulnerability detection that applies a courtroom-style paradigm.
-VulTrial comprises of four specific-role agents with distinct roles: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury).
+8. Conclusion and Future Work
+We proposed VulTrial, a novel explainable multi-agent LLM-based approach for vulnerability detection that applies a courtroom-style paradigm.
+VulTrial comprises four role-specific agents: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury).
 Each agent is essential to get the best performance (Section 4.4).
-We evaluated VulTrial against single-agent (Ding et al.’s CoT) (Ding et al., 2025b) and multi-agent, GPTLens (Hu et al., 2023), vulnerability detection baselines.
-Our results show that VulTrial outperforms both the single-agent and multi-agent baselines.
-In experiments using GPT-3.5, VulTrial enhances the performance of the single-agent baseline by 151.69%, and outperforms GPTLens by 239.78%.
-Additionally, instruction tuning with just 50 sample pairs yields further performance improvement.
-When utilizing GPT-4o with instruction tuning in moderator agent, VulTrial improves single-agent and multi-agent baselines by 139.89% and 118.3%.
-Furthermore, the number of discussion rounds among the first three agents in VulTrial does not improve the overall performance.
-For future direction, we plan to apply VulTrial to different domains to evaluate its generalizability further.
-We also plan to test additional LLMs such as DeepSeek-R1 (DeepSeek-AI et al., 2025) and LLaMA3 (Grattafiori et al., 2024) and consider each model’s strengths and weaknesses to choose the one that best fits for each role-specific task.
+We evaluated VulTrial against single-agent (CodeBERT (feng2020codebert), CodeT5 (wang2021codet5), UniXCoder (guo2022unixcoder), LineVul (fu2022linevul), and Ding et al.’s CoT (ding2024vulnerability)) and multi-agent (GPTLens (hu2023large)) baselines.
+Our results show that VulTrial outperforms both the single-agent and multi-agent baselines, with instruction tuning in the moderator agent further enhancing its performance.
+We have also demonstrated the efficacy of VulTrial across different LLMs, including an open-source, in-house-deployable model (LLaMA-3.1-8B), as well as the high quality of its generated explanations and its ability to uncover multiple zero-day vulnerabilities.
+For future work, we plan to apply VulTrial to more datasets to evaluate its generalizability further.
+We also plan to consider the strengths and weaknesses of different LLMs to choose the one that best fits as the base model for each role-specific task.
+Acknowledgements.
+This research / project is supported by the National Research Foundation, Singapore, and the Smart Nation Group under the Smart Nation Group’s Translational R&D Grant (Award No.
+TRANS2023-TGC02).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore or the Smart Nation Group.